### v3

# Evan Munro    Xu Kuang    Stefan Wager

Abstract
Policy-relevant treatment effect estimation in a marketplace setting requires taking into account both the direct benefit of the treatment and any spillovers induced by changes to the market equilibrium. The standard way to address these challenges is to evaluate interventions via cluster-randomized experiments, where each cluster corresponds to an isolated market. This approach, however, cannot be used when we only have access to a single market (or a small number of markets). Here, we show how to identify and estimate policy-relevant treatment effects using a unit-level randomized trial run within a single large market. A standard Bernoulli-randomized trial allows consistent estimation of direct effects, and of treatment heterogeneity measures that can be used for welfare-improving targeting. Estimating spillovers—as well as providing confidence intervals for the direct effect—requires estimates of price elasticities, which we provide using an augmented experimental design. Our results rely on all spillovers being mediated via the (observed) prices of a finite number of traded goods, and the market power of any single unit decaying as the market gets large. We illustrate our results using a simulation calibrated to a conditional cash transfer experiment in the Philippines.
Keywords: Equilibrium Effects, Experimental Design, Interference
1 Introduction
The standard analysis of treatment effects in a randomized trial relies on an assumption that the treatment status of one individual does not affect the outcomes of other individuals (Fisher 1935, Banerjee & Duflo 2011, Imbens & Rubin 2015). When an intervention affects supply or demand and market prices affect the outcome of interest, then general equilibrium effects lead to spillover effects, and typical estimators using data from randomized trials do not correspond to a meaningful measure of policy impact (Heckman et al. 1998).
In practice, the most common approach to address interference of this type is clustering units at a higher-level at which there are no spillover effects (Baird et al. 2018, Hudgens & Halloran 2008). Examples include the evaluation of general equilibrium impacts of cash transfers in Banerjee et al. (2021), Egger et al. (2022) and Filmer et al. (2023), and incentives for suppliers or consumers in online two-sided markets (Brennan et al. 2022, Holtz et al. 2024). However, if a market is tightly integrated, splitting the market into effectively isolated sub-markets may not be feasible. This highlights the need for spillover-aware methods for treatment effect estimation that can be used on data from an experiment run on a single interconnected market.
In this paper, we introduce a framework for treatment effect estimation in a market equilibrium setting without relying on parametric, distributional, or homogeneity assumptions. Rather than relying on market splitting, we leverage a structural assumption whereby all spillovers are mediated through the (observed) equilibrium prices of a finite number of traded goods. Our main finding is that integrating such price equilibrium effects into the potential-outcomes model for causal inference enables us to estimate policy-relevant treatment effects using unit-level experiments. In particular, we show how our approach yields asymptotic characterizations of policy-relevant average and conditional average treatment effects, and motivates the design of new forms of randomized trials and targeting rules.
We frame our model in terms of a potential outcomes specification that allows for cross-unit interference (Hudgens & Halloran 2008, Manski 2013, Aronow & Samii 2017). When equilibrium effects are removed, the model reduces to the standard Neyman-Rubin potential outcomes framework (Imbens & Rubin 2015). In a market with $n$ participants, an intervention is assigned randomly conditional on pre-treatment covariates. Each individual’s production and consumption choices are determined by latent supply and demand functions that may be shifted by the intervention (Angrist et al. 2000, Heckman & Vytlacil 2005). Units interact via a market modeled as a multiple goods economy in general equilibrium, where individuals’ production and consumption decisions depend only on their own type, their own treatment, and their expectation of the market price (Walras 1900, Mas-Colell et al. 1995). In this setup, equilibrium prices induce an exposure mapping in the sense of Manski (2013) and Aronow & Samii (2017): One unit’s treatment impacts another’s outcomes only through the treatment’s impact on the price equilibrium, which rules out peer effects or other forms of network-type interference. The price ensures that average net demand (demand minus supply) is equal to zero. The validity of this restriction on spillovers is context-specific, and depends both on the nature of the intervention and the type of market.
We start our analysis by considering the marginal effect of increasing treatment probabilities for all market participants; following Carneiro et al. (2010), we refer to this quantity as a marginal policy effect. As shown in Hu et al. (2022), the marginal policy effect can be disposed into a direct and indirect treatment effect, where the indirect effect is a general measure of spillover effects. We then analyze the sample average direct and indirect effects of a binary treatment in equilibrium (Halloran & Struchiner 1995, Hu et al. 2022, Sävje et al. 2021). In finite samples, these estimands are difficult to characterize statistically because the price equilibrium induces a complex form of data dependence. We show, however, that under an assumption that the $n$ participants in the market are independently sampled from a population distribution, we can use techniques based on empirical-process theory to derive simple large-sample limiting expressions for the direct and indirect effects as $n$ grows. These limiting expressions can then be used to guide estimation and inference for both sample and population treatment effects.
In our setting, the average direct effect can be consistently estimated via a simple difference in means estimator that ignores spillovers. Consistent estimation of the average indirect effect—and construction of confidence intervals for either estimand—also requires estimates of price elasticities. We consider augmenting standard Bernoulli-randomized experiments with small, random price perturbations that do not disrupt the overall market equilibrium. The use of such random price perturbations is in line with the documented practice of online marketplaces using randomly assigned coupons, discounts or bonuses to learn about price elasticities (Castillo 2023). We then show how such augmented experiments can be used to provide valid large-sample inference for both the average direct and indirect effects using data from a single market.
Our model allows for unrestricted heterogeneity across agents, and thus also enables us to estimate how treatment effects vary with pre-treatment characteristics—and to exploit this heterogeneity to learn improved policies that can be deployed in a market similar to the one observed. We propose versions of the sample direct effect and indirect effect that are conditional on pre-treatment covariates. These estimands depend both on the population distribution and the size of the market. They are related to the large literature on targeting without spillovers (Manski 2004, Athey & Imbens 2016, Wager & Athey 2018, VanderWeele et al. 2019), in that the conditional average direct effect is equal to the conditional average treatment effect when interference is removed. We show that as the market size grows large, the conditional estimands converge to population estimands that are simple combinations of price elasticities and conditional expectations of the direct effect of the treatment on outcomes and net demand. Furthermore, we show that it is possible to use data from a standard randomized experiment (without price perturbations) to estimate the outcome-maximizing treatment rule within the class of all treatment rules that do not move the observed market equilibrium. This optimal rule takes the form of a linear thresholding rule in the space of conditional average direct effects on outcomes and net demand.
Filmer et al. (2023) use a village-level cluster randomized experiment to demonstrate that the Pantawid conditional cash transfer program in the Philippines affects the equilibrium price of perishable protein sources, such as eggs, which leads to spillover effects on health outcomes for children in non-eligible families. In Section 5, we argue that this application fits the assumptions of our paper. We then use the data from this setting to estimate an equilibrium model of children’s health outcomes and the market for eggs in a small village in the Philippines, and demonstrate the consistency and coverage properties of our treatment effect and variance estimators based on an augmented individual-level randomized experiment generated by simulating data from this model. We also augment the model to include heterogeneous responses to the cash transfer and describe the properties of the equilibrium-stable targeting rule on samples drawn from this model.
1.1 Related Work
Much of the existing work on treatment effect estimation with spillovers focuses on setting where the Stable Unit Treatment Value Assumption (SUTVA) holds for higher-level clusters of units (Baird et al. 2018, Basse et al. 2019, Hudgens & Halloran 2008, Karrer et al. 2021, Liu & Hudgens 2014, Tchetgen & VanderWeele 2012). Another related approach is the network interference model, which posits that interference operates along a network and the connections between units are sparse (Athey et al. 2018, Leung 2020, Sävje et al. 2021). The sparsity of connections then enables randomization-based methods for studying cluster-randomized experiments to be extended to this setting. In our setting, however, the interference pattern produced by marketplace price effects is dense and simultaneously affects all units, so neither cluster nor sparsity-based methods are applicable.
Our use of a random sampling model presents a departure from the literature cited above, where typically the sample is held fixed and inference is only driven by random treatment assignment. This random sampling model, however, plays a key role in enabling our analysis, as it allows us to leverage tools from empirical process theory (e.g, van der Vaart & Wellner 1996) to address difficulties around potentially unstable and non-unique equilibrium formation in finite samples. [^1] As discussed in detail in the following section, we assume that there is a unique market-clearing price in the population; however, we allow each market participant to respond discontinuously to market prices and this may lead to non-uniqueness of possible equilibria in finite samples. In doing so, our paper adds to a handful of recent papers that use a stochastic model in a causal inference setting. Johari et al. (2020) use a mean-field model to examine the benefits of a two-sided randomization design in two-sided market platforms and quantify the bias due to interference between market participants. Li & Wager (2022) study large sample treatment effect estimation under network interference, where the exposure graph is randomly generated from a graphon model. Finally, Wager & Xu (2021) consider a model where exogenous demand is matched with endogenous supply, with interference via supply cannibalization. They propose using zero-mean perturbations to estimate a revenue gradient and optimize a continuous decision variable dynamically. Our work also uses zero-mean perturbations as part of a unit-level experiment, but is otherwise distinct, given our focus is on treatment effects of a binary intervention in a general price equilibrium.
Our model combines a potential outcomes model with some limited structural assumptions on the price equilibrium. The insight that structural modeling can be used to address endogeneity issues in causal inference has a long tradition in economics (Heckman 1979, Roy 1951, Heckman et al. 1998), and there is a large literature that uses structural modeling to develop econometric methods for causal inference in models with endogenous treatment selection (Angrist et al. 2000, Carneiro et al. 2011, Heckman & Vytlacil 2005). To date, however, there has been less emphasis on methods for estimating equilibrium effects in systems with randomized trials, where treatment selection is not an issue.
We do note a growing literature that combines data from a randomized trials with parametric structural models to estimate both partial and general equilibrium policy effects. For example, Neilson et al. (2019) uses a randomized control trial to estimate the direct effect of information provision on school choice but simulates a fully parametric structural model to evaluate equilibrium effects of the treatment. In contrast, our approach uses a more complex unit-level experiment and the structure of the price equilibrium to analyze equilibrium effects without imposing a parametric model. To accomplish this, we rely on a representation of population estimands as a combination of price elasticities and expectations of the outcome and demand distribution. This is related to the sufficient statistics approach, as described in Chetty (2009), which also expresses policy effects in terms of combinations of estimable elasticities without imposing a specific parametric model, although this is usually in a partial equilibrium setting. In the macroeconomics literature, Wolf (2023) uses a semi-structural approach to decompose the effect of economic shocks into partial equilibrium effects and general equilibrium effects (e.g., price effects), but uses time series methods to estimate the price effects, in contrast to the unit-level approach presented in this paper.
The literature on targeting treatments under interference is quite limited. Viviano (2019) estimates treatment allocation rules in a network interference model, where his results depend on sparsity of the network. Sahoo & Wager (2022) study policy learning with strategic agents, where the policymaker has a budget constraint, which leads to a type of interference. Both approaches rely on a form of direct empirical welfare maximization and do not define spillover-aware measures of treatment heterogeneity as we do in this paper.
Finally, the price perturbations that we use to estimate average price elasticities can be thought of as generating an ideal instrument for our setting. By construction, everyone in the sample is a complier, and the size of the perturbations shrink so that local elasticities can be estimated without imposing parametric assumptions on demand or supply. In empirical work with observational data, instrumental variables approaches have been used in both parametric and non-parametric settings to estimate price elasticities or a weighted average of elasticities (Angrist et al. 1996, Berry & Haile 2021). Angrist & Krueger (2001) reviews IV approaches and their relation to causal inference for observational data. In settings where it is not possible to randomize individual-level fees, under additional structural assumptions, a traditional IV-based approach could be used to estimate relevant elasticities from observational data.
2 Model and Estimands
We work under a potential-outcomes model of a market with $n$ participants, where each participant is drawn independently from a population. A binary intervention affects individual supply, demand and outcomes in the market. We use this model to define the estimands of interest for our paper, both at the sample level and at the population level. We use “sample” to refer to a market with $J$ goods and $n$ participants, since we assume that the observed equilibrium arises in this market and in many cases we will observe outcomes for all $n$ market participants. However, as we explain in more detail in Section 3, our inference approach will also allow for settings where we can only observe outcomes for a sub-sample of the $n$-sized market (but equilibrium effects are still determined by the full size-$n$ sample).
In a market of size $n$, each market participant $i=1,\,\ldots,\,n$ is characterized by an outcome function $Y_{i}(w_{i},\,p)$, which is a map from $\{0,1\}\times\mathbb{R}^{J}\to\mathbb{R}$, a net demand function $Z_{i}(w_{i},\ p)$, which is a map from $\{0,1\}\times\mathbb{R}^{J}\to\mathbb{R}^{J}$, and covariates $X_{i}\in\mathcal{X}$. These quantities are drawn jointly from a population distribution. Net demand is defined as demand minus supply. We also allow for an exogenous individual-level price shifter $U_{i}\in\mathbb{R}^{J}$. For our analysis of classical randomized trials, $U_{i}=0$, but we will also consider a class of experiments that randomize discounts or fees in a market, where $U_{i}\neq 0$ and is also exogenous. Each market participant is also assigned a treatment $W_{i}\in\{0,1\}$. Using these variables, we can now define potential outcomes in our model. Throughout this paper, we use bold-face notation to denote a vector-valued quantity collected across all market participants; for example, we write $\mathbf{Z}=(Z_{1},\ \ldots,\,Z_{n})$.
Assumption 1.
[…]
$\mathcal{S}_{\mathbf{w}}=\left\{p\in\mathbb{R}^{J}:\left\lVert\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(w_{i},p+U_{i})\right\rVert_{2}\leq a_{n}\right\}$ (1)
is non-empty with probability at least $1-e^{-cn}$ for all $n$. On the event where this set is non-empty, the market price is in this set, $P_{n}(\mathbf{w})\in\mathcal{S}_{\mathbf{w}}$.
In Section 3, we will use observations of net demand, outcomes, treatments and price-shifters in a single market of size $n$ to estimate policy-relevant treatment effects. Assumption 1 is the main structural assumption in the paper that makes this possible. Treatment and market prices directly impact outcomes and net demand for each individual. Prices are endogenous, and are determined by a sample equilibrium condition on net demand. Because of this endogeneity, there are spillovers from individual $i$ to individual $j$ through their effect on the equilibrium that the market reaches. Although our model allows for equilibrium spillovers, it rules out other types of spillovers, such as through peer effects or some other network mechanism. The validity of this restriction on interference primarily depends on the nature of the intervention and the context of the market. A subsidy allocated to suppliers in an online market may, by design, be impossible to share with those in the control group, so spillovers occur only through market prices. On the other hand, an information intervention that affects consumption in a village economy may have spillovers both through social networks in the village and through the market equilibrium.
The second part of the assumption characterizes how prices form in finite samples. The finite market price does not need to exactly clear the market; instead, we impose a restriction on the rate at which the error in the market-clearing condition shrinks with the sample size. Furthermore, it does not need to be the unique price meeting this condition, although we will require uniqueness at the population level, i.e., in the limit as the sample size grows large (Assumption 2). Proposition 21 in Appendix D provides explicit sufficient conditions for Assumption 1 to hold in a single good market without price shifters and with discontinuous net demand functions.
For a market with $n$ participants where a vector of treatments $\bm{W}$ and matrix of price perturbations $\mathbf{U}$ is realized, we observe $Y_{i}=Y_{i}(\bm{W})$ and we also observe realized net demand $Z_{i}=Z_{i}(W_{i},\bm{P}_{n}(\bm{W})+U_{i})$. For the majority of the paper, we will consider settings without price shifters (i.e., with $\mathbf{U}=0$), where treatments are Bernoulli-randomized in the sample as described below. In Section 3, we will also consider a class of augmented randomized trials with exogenous price shifters $U_{i}$ whose size shrinks as the market size grows large (Design 2).
Design 1.
Bernoulli Randomized Trial. Each unit has a known randomization probability $\pi_{i}$ satisfying $\eta\leq\pi_{i}\leq 1-\eta$ for some $\eta>0$, such that treatment is generated as $W_{i}\sim\text{Bernoulli}(\pi_{i})$. The $\pi_{i}$ may be constant, functions of $X_{i}$, or randomly generated. There are no price perturbations, so $U_{i}=0$ for all $i\in\{1,\,\ldots,n\}$.
Given this causal model, there are many treatment effects one can seek to study. The total treatment effect is the average difference in outcomes from treating everyone versus no one in the market,
$\tau_{\text{TOT}}=\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,P_{n}(\mathbf{1}_{n})+U_{i})-Y_{i}(0,\,P_{n}(\mathbf{0}_{n})+U_{i})\right),$ (2)
where $\mathbf{1}_{n}$ and $\mathbf{0}_{n}$ are $n$-length vectors of 1s and 0s. Total treatment effects can readily be estimated across disconnected markets using cluster-randomized experiments clustered at the market level (Baird et al. 2018). However, without further assumptions, estimating $\tau_{\text{TOT}}$ from a single (connected) market is impossible, as it requires predicting behavior at prices that are far from anything seen in the experiment. A more tractable—and still fruitful—approach is to focus on treatment effects that, in the sense of Carneiro et al. (2010), are marginal to the treatment-assignment policy used to collect the data. For any $n$-length vector of Bernoulli randomization probabilities $\bm{\pi}$, write
$V_{n}(\bm{\pi})=\mathbb{E}_{\bm{\pi}}\left[\frac{1}{n}\sum_{i=1}^{n}Y_{i}\right],$ (3)
where we use the notation $\mathbb{E}_{\bm{\pi}}\left[\cdot\right]=\mathbb{E}\left[\cdot\,\big{|}\,\left\{Y_{i}(w,\,p),\,Z_{i}(w,\,p),U_{i}\right\}_{i=1}^{n}\right]$ to denote expectations over the Bernoulli treatment assignment while holding the sample (i.e., the potential outcomes) fixed. The marginal policy effect in the sample, $\tau_{\text{MPE}}$, is a local approximation to the total treatment effect defined in (2). It measures the average effect of increasing treatment probabilities for everyone:
$\tau_{\text{MPE}}=\frac{d}{d\varepsilon}V(\bm{\pi}+\bm{\varepsilon})|_{\varepsilon=0}.$ (4)
Hu et al. (2022) show that the marginal policy effect can be decomposed into the sum of two sample average treatment effects, which are the first set of target estimands in our paper. Here, the notation $(W_{i}=a;\bm{W}_{-i})$ represents the random vector we obtain by first sampling $\bm{W}$, and then setting its $i$-th coordinate to a fixed value $a$:
$\begin{split}&\tau_{\text{MPE}}=\tau_{\text{ADE}}+\tau_{\text{AIE}},\\ &\tau_{\text{ADE}}=\frac{1}{n}\sum\limits_{i=1}^{n}\mathbb{E}_{\pi}[Y_{i}(W_{i}=1,P_{n}(W_{i}=1;\bm{W}_{-i})+U_{i})-Y_{i}(W_{i}=0,P_{n}(W_{i}=0;\bm{W}_{-i})+U_{i})],\\ &\tau_{\text{AIE}}=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{j\neq i}\mathbb{E}_{\pi}[Y_{j}(W_{j},P_{n}(W_{i}=1;\bm{W}_{-i})+U_{i})-Y_{j}(W_{j},P_{n}(W_{i}=0;\bm{W}_{-i})+U_{i})].\end{split}$ (5)
$\tau_{\text{ADE}}$ is the average effect on an individual’s outcome of changing their own treatment in a market with $n$ participants, and is a standard quantity in the causal inference literature (e.g., Halloran & Struchiner 1995, Sävje et al. 2021). It is a measure of the direct effect of the treatment while holding the equilibrium fixed. When SUTVA holds, then $\tau_{\text{ADE}}$ is equal to the sample Average Treatment Effect (SATE) and $\tau_{\text{AIE}}$ is 0. The average indirect effect in the sample is the average effect on everyone else’s outcomes of changing an individual’s treatment. It is a general measure of spillover effects.
The estimands $\tau_{\text{ADE}}$ and $\tau_{\text{AIE}}$ are defined at the sample level. We also find it useful to introduce population versions of our target estimands, and define the following whenever the limits below exist:
$\begin{split}&\tau^{*}_{\text{ADE}}=\lim\limits_{n\rightarrow\infty}\mathbb{E}[\tau_{\text{ADE}}]=\lim\limits_{n\rightarrow\infty}\mathbb{E}[Y_{i}(W_{i}=1;\bm{W}_{-i})-Y_{i}(W_{i}=0;\bm{W}_{-i})],\\ &\tau^{*}_{\text{AIE}}=\lim\limits_{n\rightarrow\infty}\mathbb{E}[\tau_{\text{AIE}}]=\lim\limits_{n\rightarrow\infty}(n-1)\cdot\mathbb{E}[Y_{j}(W_{i}=1;\bm{W}_{-i})-Y_{j}(W_{i}=0;\bm{W}_{-i})].\end{split}$ (6)
In the next section, we will show that under our model and an additional set of regularity conditions, these limits in fact exist and the population estimands have a tractable representations. These representations will then be useful for designing estimators that are consistent for both sample and population-level targets. In addition, just as in the literature on causal inference under SUTVA, in some settings the population estimand is the more relevant target. For example, we may have data from a small market and are interested in doing inference on a similar but larger market, in a setting where there is significant homogeneity across markets.
2.1 Large Sample Characterization
Our asymptotic characterization of the average direct and indirect effect requires some additional regularity conditions. It is first helpful to introduce some additional notation for moments of the population distribution. Let $y(p)=\mathbb{E}[Y_{i}(W_{i},p)]$ and $z(p)=\mathbb{E}[Z_{i}(W_{i},p)]$. For $w\in\{0,1\}$, let $y(w,p)=\mathbb{E}[Y_{i}(w,p)]$ and $z(w,p)=\mathbb{E}[Z_{i}(w,p)]$. Define $p^{*}_{\pi}$ to be the price that clears the population market when the treatment is allocated according to $\pi(\cdot)$, i.e., $p^{*}_{\pi}=\{p:z(p)=0\}$.
We start by making regularity assumptions on the expected net demand, and in particular assume that there is a unique equilibrium price in the population. Market values will be in a compact set as long as all net demand functions are weakly negative for prices above an upper bound and weakly positive for prices smaller than some lower bound. In a single good market, uniqueness of $p^{*}_{\pi}$ only requires continuity and strict monotonicity of $z(p)$ in $p$. In a market with multiple goods, uniqueness and existence of the equilibrium price can be shown by ensuring the expected net demand function is a contraction, as is common in the literature on models of strategic behavior (Cornes et al. 1999, Van Long & Soubeyran 2000). Other approaches that rely on more primitive assumptions are also possible; under a zero-degree homogeneity assumption in net demand then a gross-substitutes condition ensures uniqueness (Arrow & Hahn 1971).
Assumption 2.
Market prices take values in a compact set $\mathcal{S}\subset\mathbb{R}^{J}$ almost surely. Given any randomization policy $\pi$, there is a unique population market-clearing price $p^{*}_{\pi}\in\mathcal{S}$ that satisfies $z(p^{*}_{\pi})=0.$ Let the Jacobian $\xi_{z}=\nabla_{p}z(p^{*}_{\pi})$, where row $j$ of the $J\times J$ matrix is $\nabla^{\top}_{p}z_{j}(p^{*}_{\pi})$. The Jacobian is full rank.
Next, we make regularity assumptions on the unit-level net demand and outcome functions. To ensure that a variety of random processes studied in the paper concentrate, Assumption 3 and Assumption 4 impose some constraints on outcomes and net demand at an individual level. These restrictions have a simple economic interpretation and are general enough to encompass a wide variety of possible data-generating processes that underlie market behavior. Assumption 3 is our main assumption on the net demand, and is a generalization of the familiar assumption in a single-good setting that net demand be non-increasing in price. In the multiple good case, we also require that every unit $i$ will decrease their net demand of good $j$ more in response to a decrease in the price of good $j$ than in response to a relatively small change in the price of other goods. This assumption restricts the magnitudes of cross-price elasticities compared to own-price elasticities.
Assumption 3.
[…]
$Z_{ij}(w,\,p-\varepsilon e_{j})\geq Z_{ij}(w,\,p+\delta)\geq Z_{ij}(w,\,p+\varepsilon e_{j}),\ \text{ for all }\ \left\lVert\delta\right\rVert_{2}\leq C\varepsilon,$ (7)
where $e_{j}$ denotes the $j$-th basis vector.
Assumption 4 allows the outcome function to be the sum of a term $H_{i}(w,\,p)$ with a Lipschitz dependence on prices, and a bilinear function of prices and net demand. Here, $H_{i}(w,\,p)$ could be a direct measurement of welfare, which under standard rational choice assumptions should vary continuously with prices. Or, it could be any Lipschitz transformation of net demand and prices when net demand is itself Lipschitz. When net demand is not smooth at the individual level, we allow for a bilinear function of $p$ and $Z_{i}(w,\,p)$, which captures natural researcher-constructed outcomes such as surplus or profit. This allows for outcomes that are discontinuous in prices, but the discontinuity must only through simple transformations of net demand. We note that all our results below also hold if $Y_{i}(w,\,p)$ depends on $\left(\alpha+\beta^{\top}p\right)G_{i}(w,\,p)$ for any function $G_{i}(w,\,p)$ satisfying Assumption 3; however, for simplicity, we do not add this degree of flexibility in the notation. [^2] This captures researcher-constructed outcomes that depend on possibly discontinuous consumption or production of certain goods in a monotonic way. For example, in Section 5, health outcomes depend on consumption of protein sources, and this relationship could be non-linear in our framework, even when individual-level consumption is discontinuous.
Assumption 4.
The outcome functions can be decomposed as the sum of a Lipschitz welfare function and a (random) function of $p$ times net demand,
$Y_{i}(w,\,p)=H_{i}(w,\,p)+\left(\alpha+\beta^{\top}p\right)\cdot Z_{i}(w,\,p),$ (8)
where $H_{i}(w,\,p)$ is $M$-Lipschitz in $p$, $\alpha\in\mathbb{R}^{J}$ and $\beta\in\mathbb{R}^{J\times J}$.
Finally, in Assumption 5 we list some additional regularity assumptions, which are standard in the literature on asymptotic statistics. Although individual net demand and outcome functions may be discontinuous, we require their expectation to vary smoothly in prices. The weak continuity assumption effectively requires that discontinuity points cannot concentrate at specific values of $p$; for example, in a one-dimensional threshold-crossing model, our weak continuity assumption is equivalent to the cutoff point distribution having a density that is uniformly bounded from above.
Assumption 5.
The following regularity conditions hold:
  1. 1.
    Net demand and outcome functions are uniformly bounded, i.e., there is a constant $M<\infty$ such that, almost surely, $\left\lvert Y_{i}(w,\,p)\right\rvert\leq M$ and $\left\lvert Z_{ij}(w,\ p)\right\rvert\leq M$ for all $w\in\{0,1\}$, $p\in\mathcal{S}$ and $j\in\{1,\ldots,J\}$.
  2. 2.
    Net demand is weakly continuous in $p$. There is a constant $L>0$ such that for all pairs of prices $p,\,p^{\prime}$, all $w$, and all $j$, we have $\mathbb{E}[(Z_{ij}(w,p)-Z_{ij}(w,p^{\prime}))^{2}]\leq L\left\lVert p-p^{\prime}\right\rVert_{2}$.
  3. 3.
    For all $p\in\mathcal{S}$ and $w\in\{0,1\}$, $y(w,p)$, $z(w,p)$, $y(p)$ and $z(p)$ are twice continuously differentiable in $p$ with bounded first and second derivatives.
  4. 4.
[…]
Our first result is that, given these assumptions, the sample equilibrium price concentrates on $p^{*}_{\pi}$ as the sample size grows, and has an asymptotically linear representation (i.e., to first order, random price fluctuations can be written as a sum of additive contributions from each unit). All proofs are given in the appendix. Our technical results rely heavily on asymptotic equicontinuity results from empirical process theory as described in van der Vaart & Wellner (1996).
Theorem 1.
Under Assumptions 1, 2, 3 and 5, the finite sample equilibrium price under Design 1 and Design 2 satisfies
$P_{n}(\bm{W})-p^{*}_{\pi}=-\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi})+o_{p}(n^{-1/2}),$ (9)
and furthermore $\sqrt{n}\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\Rightarrow\mathcal{N}\left(0,\,\xi_{z}^{-1}\operatorname{Var}\left[Z_{i}(W_{i},p^{*}_{\pi})\right]\left(\xi_{z}^{-1}\right)^{\top}\right)$, where $\xi_{z}$ is as defined in Assumption 2.
This result is a crucial building block towards the rest of the theory in the paper, since it provides a representation of $P_{n}(\bm{W})$ in terms of the fixed price $p^{*}_{\pi}$ and an average of IID terms. This will be helpful in characterizing the asymptotic variance of various random variables in our model. Before introducing our estimators in Section 3, we next apply this result to characterizing the relationship between the sample and population-level average direct and indirect effect. To this end, we first show that $\tau^{*}_{\text{ADE}}$ and $\tau^{*}_{\text{AIE}}$ exist and have a simple representation in terms of moments of the population distribution.
Theorem 2.
Under Assumptions 1, 2, 3, 4 and 5, the population-level estimands in the Bernoulli Design 1 are:
$\begin{split}&\tau^{*}_{\text{ADE}}=y(1,p^{*}_{\pi})-y(0,p^{*}_{\pi}),\\ &\tau^{*}_{\text{AIE}}=-\xi_{y}^{\top}\xi_{z}^{-1}[z(1,p^{*}_{\pi})-z(0,p^{*}_{\pi})],\end{split}$ (10)
where $\xi_{y}=\nabla_{p}\mathbb{E}[Y_{i}(W_{i},p^{*}_{\pi})]$ is a $J\times 1$ vector and $\xi_{z}$ is as defined in Assumption 2.
The population direct effect is the difference in expected outcomes at $p^{*}_{\pi}$ evaluated at $W_{i}=1$ and $W_{i}=0$. In the next section, we show that this can be estimated with a differences in means estimate. This finding is in line with Sävje et al. (2021), who show that estimators that target the average treatment effect under no-interference settings generally recover the average direct effect under interference. However, our results will go beyond those of Sävje et al. (2021), since we will also provide a central limit theorem for estimators of $\tau^{*}_{\text{ADE}}$.
The indirect effect, which is a key component of the policy counterfactual $\tau_{\text{MPE}}$, is not estimable using variation in treatment only. The population indirect effect depends on the product of three terms: how the treatment affects net demand, how net demand is affected by prices ($\xi_{z}$), and how changes in prices affect outcomes ($\xi_{y}$). In settings where the treatment impacts market-clearing prices through net demand, and outcomes are sensitive to market prices, then spillover effects and the indirect effect are stronger. In Section 3 we will show how each component of $\tau^{*}_{\text{ADE}}$ and $\tau^{*}_{\text{AIE}}$ can be estimated using unit-level experiments.
Finally, we connect the limiting estimands $\tau^{*}_{\text{ADE}}$ and $\tau^{*}_{\text{AIE}}$ to their finite-sample counterparts $\tau_{\text{ADE}}$ and $\tau_{\text{AIE}}$. In the case of the direct effect, Theorem 3 provides an asymptotically linear expansion for $\tau_{\text{ADE}}$ around $\tau^{*}_{\text{ADE}}$, thus allowing us to disambiguate between the sample and population estimands when conducting inference. This mirrors the well-known relationship between the sample- and population-average treatment effects in the no-interference setting (Imbens 2004).
Theorem 3.
Under the conditions of Theorem 2,
$\tau_{\text{ADE}}=\frac{1}{n}\sum\limits_{i=1}^{n}\Big{(}Y_{i}(1,p^{*}_{\pi})-Y_{i}(0,p^{*}_{\pi})-\left(\pi_{i}\Delta_{i}(1,\,p^{*}_{\pi})+(1-\pi_{i})\Delta_{i}(1,\,p^{*}_{\pi})\right)\Big{)}+o_{p}(1/\sqrt{n}),$ (11)
where
$\Delta_{i}(w,\,p)=\nabla_{p}^{\top}[y(1,p)-y(0,p)]\xi_{z}^{-1}Z_{i}(w,\,p).$ (12)
Furthermore, writing $\varepsilon_{i}(w)=Y_{i}(w,p^{*}_{\pi})-y(w,p^{*}_{\pi})$, we have
$\sqrt{n}\left(\tau_{\text{ADE}}-\tau^{*}_{\text{ADE}}\right)\Rightarrow\mathcal{N}\left(0,\,\operatorname{Var}\left[\varepsilon_{i}(1)-\varepsilon_{i}(0)-\left(\pi_{i}\Delta_{i}(1,\,p^{*}_{\pi})+(1-\pi_{i})\Delta_{i}(1,\,p^{*}_{\pi})\right)\right]\right).$ (13)
Relative to the no-interference setting, the expansion in Theorem 3 includes additional $\Delta_{i}$ terms that depend on individual-level net demand. In a finite market where treatment affects demand and supply, there is variation in realized market prices that then impacts $\tau_{\text{ADE}}$. The $\Delta_{i}$ terms capture this additional source of variation. We also note that, as in the no-interference setting, the variance of $\tau_{\text{ADE}}$ is generally not identified, as it depends on the covariance of individual treated and control potential outcomes; however, as shown in Proposition 6, it will still be possible to provide asymptotically conservative inference (i.e., construct confidence intervals with potentially greater-than-nominal coverage).
Meanwhile, for the indirect effect, $\tau_{\text{AIE}}$ converges to $\tau_{\text{AIE}}^{*}$, but Theorem 4 does not provide a rate of convergence. We conjecture that establishing useful rates of convergence here would require replacing Assumptions 2 and 3 with stronger assumptions that are more explicit about finite-sample price formation. Under our current assumptions, the convergence rate of $\tau_{\text{AIE}}$ can be slower than $\sqrt{n}$; see Appendix D.2 for additional explanation in the context of a simple example. Since our estimator in Section 3 targeting $\tau_{\text{AIE}}^{*}$ also converges at a slower than $\sqrt{n}$ rate, we conjecture that confidence intervals built for $\tau^{*}_{\text{AIE}}$ will also generally cover $\tau_{\text{AIE}}$; we also note that this holds in our numerical experiments.
Theorem 4.
[…]
3 Estimation and Inference
We now move to estimation of the quantities defined in the previous section using data from unit-level randomized experiments. We first derive the the limiting distribution of a differences-in-means estimator when data is generated from a standard RCT that is run on the entire market. This estimator is consistent for the average direct effect, but its asymptotic variance depends on price-elasticity terms that cannot be estimated using data from an RCT that randomizes treatment only. We show that if confidence intervals are constructed using a variance estimator based on the asymptotic variance without price interference, then coverage for the population direct effect will not be asymptotically exact.
In order to perform inference for the direct effect and to perform estimation and inference for the indirect effect, we introduce an augmented randomized trial, with Bernoulli randomization for the treatment and local perturbation of market prices to estimate price elasticities. The estimator for the direct effect remains unchanged from the standard RCT. Algorithmically, the estimator for the indirect effect looks algorithmically like a combination of differences-in-means estimator and an instrumental variables estimator. Under the augmented randomized experiment, we show how to construct asymptotically valid confidence intervals for population versions of both effects. In this section, we assume that experiments are run on the entire market of $n$ participants. However, as described in more detail in Appendix D.3, the extension to settings where only a sub-sample of the market is observed is straightforward.
3.1 Estimation for the Direct Effect in a Standard RCT
Following a number of recent papers, including Sävje et al. (2021) and Li & Wager (2022), we consider estimating the direct effect using a differences-in-means estimator
$\hat{\tau}_{\text{ADE}}=\frac{1}{n}\sum\limits_{i=1}^{n}\left[\frac{W_{i}Y_{i}}{\hat{\pi}}-\frac{(1-W_{i})Y_{i}}{1-\hat{\pi}}\right],\ \ \ \ \hat{\pi}=\frac{1}{n}\sum\limits_{i=1}^{n}W_{i}.$ (14)
Our first result is that—even under interference that occurs through market prices—the differences-in-means estimator converges to the population direct effect at a $\sqrt{n}$ rate and is asymptotically normal. The variance of the difference-in-means estimator depends both on the variance of the potential outcomes—as in a standard no-interference setting—as well as an additional term ($\Delta_{i}(w,p^{*}_{\pi})$) that is a function of net demand and price sensitivity of outcomes and net demand. [^3] Our finding that the difference in means is consistent for the ADE mirrors general findings in Sävje et al. (2021); however, the fact that we get a $\sqrt{n}$ rate of convergence and a central limit theorem depends on our marketplace model.
Theorem 5.
Under the conditions of Theorem 2 and with uniform randomization probabilities $\pi_{i}=\pi$ for all $i$, when treatment is allocated according to Design 1,
$\hat{\tau}_{\text{ADE}}=\tau^{*}_{\text{ADE}}+\frac{1}{n}\sum\limits_{i=1}^{n}\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right)+o_{p}(1),$ (15)
where $\varepsilon_{i}(w)=Y_{i}(w,p^{*}_{\pi})-y(w,p^{*}_{\pi})$ and $\Delta_{i}(w,\,p)$ is as in Theorem 3. Furthermore,
$\sqrt{n}\left(\hat{\tau}_{\text{ADE}}-\tau^{*}_{\text{ADE}}\right)\Rightarrow\mathcal{N}(0,\sigma^{2}_{D}),\ \ \ \ \sigma^{2}_{D}=\mathbb{E}\left[\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},p^{*}_{\pi})\right)^{2}\right].$ (16)
Although we have shown that the familiar difference-in-means is asymptotically normal around the ADE, the above result also implies that the usual asymptotic variance derived under the no-interference setting does not match $\sigma^{2}_{D}$ in our market equilibrium model. Instead, to construct confidence intervals that are asymptotically exact for the population direct effect, we need estimates of the price elasticities that appear in the $\Delta_{i}(w,p^{*}_{\pi})$ terms. In the next section, we will develop an augmented randomized experiment that allows us to estimate these price sensitivities as well as the AIE, which is not estimable with treatment randomization only.
Finally, we note that Theorem 5 quantifies the errors of $\hat{\tau}_{\text{ADE}}$ as an estimator of the population estimand $\tau^{*}_{\text{ADE}}$. In some settings, however, it may also be of interest to provide inference about the sample direct effect $\tau_{\text{ADE}}$. Combining Theorems 3 and 5, we can verify that
$\hat{\tau}_{\text{ADE}}-\tau_{\text{ADE}}=\frac{1}{n}\sum\limits_{i=1}^{n}(W_{i}-\pi)\left(\frac{\varepsilon_{i}(1)}{\pi}+\frac{\varepsilon_{i}(0)}{1-\pi}-\left(\Delta_{i}(1,p^{*}_{\pi})-\Delta_{i}(0,p^{*}_{\pi})\right)\right)+o_{p}\left(\frac{1}{\sqrt{n}}\right),$ (17)
and furthermore that
$\begin{split}&\sqrt{n}\left(\hat{\tau}_{\text{ADE}}-\tau_{\text{ADE}}\right)\Rightarrow\mathcal{N}\left(0,\,\bar{\sigma}_{D}^{2}\right),\\ &\bar{\sigma}_{D}^{2}=\pi\left(1-\pi\right)\mathbb{E}\left[\left(\frac{\varepsilon_{i}(1)}{\pi}+\frac{\varepsilon_{i}(0)}{1-\pi}-\left(\Delta_{i}(1,p^{*}_{\pi})-\Delta_{i}(0,p^{*}_{\pi})\right)\right)^{2}\right].\end{split}$ (18)
The asymptotic variance here still depends on the market equilibrium effects; furthermore, it depends on the correlation of $\varepsilon_{i}(1)$ and $\varepsilon_{i}(0)$ and so is generally not identified. However, the following result shows that, in an extension of the classic result of Neyman (1923), $\sigma_{D}$ is a conservative upper bound for $\bar{\sigma}_{D}$ and so confidence intervals built using Theorem 5 that are exact for $\tau^{*}_{\text{ADE}}$ will also be conservative for $\tau_{\text{ADE}}$.
Proposition 6.
Under the conditions of Theorem 5, $\bar{\sigma}^{2}_{D}\leq\sigma^{2}_{D}$.
3.2 Augmented Randomized Experiment
In order to estimate spillover effects or the build confidence intervals for the average direct effect, we need to estimate price elasticities. Here, we do so by considering an augmented experiment that introduces small, unit level price perturbation into our randomization design. The net demand and outcome processes remain the same as in the previous section, but prices are shifted at an individual level by $U_{i}$ for individual $i$.
Design 2.
Augmented Randomized Trial. Treatment is assigned based on a uniform randomization probability, so $W_{i}\sim\text{Bernoulli}(\pi)$, where there exists $\eta>0$ such that $\eta\leq\pi\leq 1-\eta$. For $i\in\{1,\,\ldots,n\}$, $j\in\{1,\,\ldots,J\}$, $U_{ij}$ is drawn independently and uniformly at random from $\{-h_{n},+h_{n}\}$, where $h_{n}=cn^{-\alpha}$ with $\frac{1}{4}<\alpha<\frac{1}{2}$ and $c$ is a finite constant.
The price perturbations introduced by this experiment can be interpreted as instruments for estimating price derivatives locally. They shift demand and supply in the neighborhood of the market-clearing price, and since they are randomized, they are independent of other variables that determine demand or outcomes. In a typical two-sided market, implementing the experiment requires introducing a small random discount, fee, or subsidy to both producers and consumers for any products where the treatment is expected to significantly affect supply or demand. Our theory requires that the size of the price perturbation decreases with sample size. For a fixed market size, as the perturbation size increases, the variance of $\hat{\tau}_{\text{AIE}}$ decreases, but large price perturbations distort production and would be likely be viewed unfavorably by various market stakeholders. The simplest guidance for choosing the size of the price perturbation is to choose a value of $c$ so that for the sample size of the experiment, the size of the price perturbation is noticeable to market participants but limited to a small percentage of the current market price (e.g. 1-2%). For markets where it is not possible for reputational or legal reasons to introduce any individual-level price variation, the components of $\tau^{*}_{\text{AIE}}$ can be estimated using price variation across products or time instead, at the cost of introducing additional structural assumptions on the environment.
The experiment affects the equilibrium in finite samples, since the price perturbations affect the market clearing condition in Assumption 1. However, the asymptotically linear expansion in Theorem 1 holds under Design 1 and Design 2, so the limiting price is the same under both designs. In a qualitative sense, this result means that the price perturbations are unlikely to meaningfully affect healthy operation of the marketplace. [^4] The fact that we only consider small, unobstrusive perturbations to the market also means we do not get to observe behaviors at prices far from the large-sample equilibrium $p^{*}_{\pi}$, and so cannot identify or consistently estimate full demand curves.
We now develop a point estimator for the indirect effect. Our proposal consists of a combination of difference-in-means estimators and estimators of price elasticities. The strategy is plug-in estimation based on the functional form for $\tau^{*}_{\text{AIE}}$ derived in Theorem 2. Let $\bm{Y}$ be the $n$-length vector of outcomes, $\bm{U}$ is the $n\times J$ matrix of price perturbations, and $\bm{Z}$ is the $n\times J$ matrix of net demand observations. The estimator is
$\hat{\tau}_{\text{AIE}}=-\hat{\gamma}^{\top}\cdot\hat{\tau}^{z}_{\text{ADE}},\ \ \ \ \hat{\gamma}=(\bm{U}^{\top}\bm{Z})^{-1}(\bm{U}^{\top}\bm{Y}),$ (19)
where $\hat{\gamma}$ is a $J\times 1$ vector that estimates $[\xi^{\top}_{z}]^{-1}\xi_{y}$. The direct effect of the treatment on net demand is estimated via a difference-in-means estimator,
$\hat{\tau}^{z}_{\text{ADE}}=\frac{1}{n}\sum\limits_{i=1}^{n}\left[\frac{W_{i}{Z}_{i}}{\hat{\pi}}-\frac{(1-W_{i}){Z}_{i}}{1-\hat{\pi}}\right].$ (20)
Our next result provides a central limit theorem for our indirect effect estimator. This estimator converges at a rate that depends on the magnitude of the price perturbations $h_{n}$, and is always slower than the $\sqrt{n}$-rate obtained for the direct effect. In deriving this result, it is helpful to note that algorithmically $\hat{\tau}_{\text{AIE}}$ is the product of a rescaled instrumental variables estimator and a differences in mean estimator. We can then use standard results on the asymptotic behavior of IV estimators to guide our analysis.
Theorem 7.
Suppose the Assumptions of Theorem 2 hold and that treatment is assigned according to Design 2. Then, the estimated average indirect effect can be written as:
$\hat{\tau}_{\text{AIE}}=\tau^{*}_{\text{AIE}}-\frac{1}{nh_{n}^{2}}\,\sum\limits_{i=1}^{n}Q_{z}^{\top}U_{i}\nu_{i}(W_{i})+o_{p}\left(\frac{1}{\sqrt{n}h_{n}}\right),$ (21)
where $\nu_{i}(W_{i})=Y_{i}(W_{i},p^{*}_{\pi})-Z_{i}(W_{i},p^{*}_{\pi})^{\top}[\xi_{z}^{-1}]^{\top}\xi_{y}$ and $Q_{z}=\xi_{z}^{-1}[\tau^{*,z}_{\text{ADE}}]$. Furthermore, this estimator satisfies a central limit theorem
$\sqrt{n}{h_{n}}\left(\hat{\tau}_{\text{AIE}}-\tau^{*}_{\text{AIE}}\right)\Rightarrow\mathcal{N}(0,\sigma^{2}_{I}),\ \ \ \ \sigma^{2}_{I}=Q_{z}^{\top}\mathbb{E}[\nu^{2}_{i}(W_{i})\,I_{J\times J}]Q_{z}$ (22)
where $I_{J\times J}$ is the $J\times J$ identity matrix.
Under general patterns of interference, inference on the indirect effect is challenging (Sävje et al. 2021, Li & Wager 2022), and consistent estimators for the variance of the indirect effect are generally not available. There are two reasons why our paper overcomes this difficulty. Although our interference pattern is dense, it is structured in that all interference happens through the market price, and the market price forms by satisfying a score condition. This structure leads to an analytical functional form for the variance of the indirect effect, as given above. Second, with data from a richer randomized experiment that includes small price perturbations, we are able to estimate each component of this variance, and the variance for the indirect effect.
We estimate $\sigma_{D}^{2}$ and $\sigma_{I}^{2}$ via natural plug-in estimators. Let $n_{w}$ denote the number of units with $W_{i}=w$. The variance of the direct effect can be estimated as
$\begin{split}&\hat{\sigma}^{2}_{D}=\frac{1}{n}\sum\limits_{i=1}^{n}\left[\frac{W_{i}\hat{\varepsilon}_{i}(1)}{\hat{\pi}}-\frac{(1-W_{i})\hat{\varepsilon}_{i}(0)}{1-\hat{\pi}}-(\hat{\xi}_{y1}-\hat{\xi}_{y0})^{\top}[\hat{\xi}_{z}]^{-1}Z_{i}\right]^{2},\\ &\hat{\varepsilon}_{i}(w)=Y_{i}-\frac{1}{n_{w}}\sum\limits_{i:W_{i}=w}Y_{i},\\ \end{split}$ (23)
where, for $w\in\{0,1\}$, $\hat{\xi}_{yw}$ is a $J\times 1$ vector estimated from regressions of $Y_{i}$ on $U_{i}$ using only observations with indices in the set $\{i:W_{i}=w\}$, and $\hat{\xi}_{z}$ is a $J\times J$ matrix which is computed via regressions of net demand $Z_{ij}$ on price perturbations $U_{i}$ for $j\in\{1,\,\ldots,J\}$ and $\hat{\xi}_{y}$ is a $J\times 1$ vector computed via a regression of $Y_{i}$ on $U_{i}$. Meanwhile, for the variance of the indirect effect, the natural plug-in estimator is [^5] In Appendix D.4, we also describe a second order adjustment to $\hat{\sigma}^{2}_{I}$ that is asymptotically negligible but yields better coverage under in our experiments when the sample size is moderate.
$\hat{\sigma}^{2}_{I}=\frac{1}{n\sqrt{h}_{n}}\sum\limits_{i=1}^{n}\left((Y_{i}-{Z}_{i}^{\top}[\hat{\xi}^{\top}_{z}]^{-1}\hat{\xi}_{y})U^{\top}_{i}\hat{\xi}_{z}^{-1}\hat{\tau}^{z}_{\text{ADE}}\right)^{2}$ (24)
These variance estimates can then be paired with our asymptotic normal approximation to build confidence intervals for the direct and indirect estimates; for example, for a confidence level of 95%, the confidence intervals are constructed as
$\hat{\tau_{D}}\pm 1.96\cdot\frac{\hat{\sigma}_{D}}{\sqrt{n}},\ \ \ \ \ \hat{\tau_{I}}\pm 1.96\cdot\frac{\hat{\sigma}_{I}}{\sqrt{n}h_{n}}.$ (25)
The following result verifies validity of these variance estimators.
Theorem 8.
[…]
4 Heterogeneous Treatment Effects
So far, we defined a potential outcomes model that captured equilibrium interference but reduced to the Neyman-Rubin model without interference. We then used this model to define average direct and indirect treatment effects under general treatment allocation rules. In the previous section, we proposed estimators for these effects that relied on data generated from randomized trials where the treatment was assigned with constant probability. In this section, we return to general treatment rules and discuss heterogeneous effects and optimal targeting when there is interference through an equilibrium statistic.
The planner controls the (potentially randomized) treatment allocation function $\pi(\cdot):\mathcal{X}\rightarrow[0,1]$ where $\pi(x)=\mathbb{P}\left[W_{i}=1\,\big{|}\,X_{i}=x\right]$. The conditional expectation functions are defined as $y(w,p,x)=\mathbb{E}[Y_{i}(w,p)\,\big{|}\,X_{i}=x]$ and $z(w,p,x)=\mathbb{E}[Z_{i}(w,p)\,\big{|}\,X_{i}=x]$. We are interested both in quantifying how relevant treatment effects vary with $x$, and how this information can be used to guide choices of $\pi(\cdot)$ that achieve better outcomes.
4.1 Definitions of Conditional Treatment Effects Under Interference
Under SUTVA, the conditional average treatment effect (CATE) is defined as $\tau(x)=\mathbb{E}[Y_{i}(1)-Y_{i}(0)|X_{i}=x]$ (Imbens 2004). Without interference, the optimal unconstrained targeting rule allocates treatments only to those with a positive CATE (Manski 2004), or potentially to those whose CATE exceeds a budget-informed threshold (Bhattacharya & Dupas 2012). Such treatment assignment rules, however, are no longer optimal under interference—and the CATE is no longer even well defined.
We begin this section by proposing two definitions of conditional estimands that play a similar role to the average direct and indirect effect, but for targeted treatments: The Conditional Average Direct Effect (CADE) and the Conditional Average Indirect Effect (CAIE). The CADE
$\bar{\tau}_{\text{CADE}}(x)=\mathbb{E}[Y_{i}(W_{i}=1;W_{-i})-Y_{i}(W_{i}=0;W_{-i})|X_{i}=x]$ (26)
is the expected effect of treating an individual with covariate value $x$ on their own outcomes in a system of $n$ individuals. The CAIE
$\bar{\tau}_{\text{CAIE}}(x)=(n-1)\mathbb{E}[Y_{j}(W_{i}=1;W_{-i})-Y_{j}(W_{i}=0;W_{-i})|X_{i}=x]$ (27)
is the expected effect of treating an individual with covariate value $x$ on everyone else’s outcomes in a system of $n$ individuals. These estimands are non-random quantities that depend both on the population distribution and on the sample size. They are policy-relevant for settings where we expect to deploy a policy in a market of similar size and composition to the observed market. When interference is removed, the CADE is equal to the CATE, and the CAIE is zero.
We next connect these definitions of heterogeneous effects to conditional marginal effects; this result is valid under general patterns of interference. In Section 2, we reported the results of Hu et al. (2022), showing the sum of $\tau_{\text{ADE}}$ and $\tau_{\text{AIE}}$ is equal to the effect on average outcomes of a marginal increase in each individual’s treatment probability. Proposition 9 extends this result to our proposed heterogeneity measures.
Proposition 9.
Let $Y_{i}(\mathbf{w})$ be potential outcome functions with an arbitrary sampling distribution, and let treatment be generated as $W_{i}\sim\text{Bernoulli}(\pi_{i})$ with treatment assignment probabilities $10<\pi_{i}<$ 1 that may be dependent on the $Y_{i}(\mathbf{w})$. Then,
$\bar{\tau}_{\text{MPE}}(x)=\bar{\tau}_{\text{CADE}}(x)+\bar{\tau}_{\text{CAIE}}(x),\ \ \ \ \bar{\tau}_{\text{MPE}}(x):=\mathbb{E}\left[\frac{\partial}{\partial\pi_{k}}\sum\limits_{i=1}^{n}\mathbb{E}_{\pi}[Y_{i}(\bm{W})]\Big{|}X_{k}=x\right].$ (28)
A policymaker may be interested in taking advantage of heterogeneous responses to treatment in the population by implementing a targeting rule, rather than a treatment rule with uniform probability. Proposition 9 shows that the sum of the CADE and the CAIE is relevant for making the decision on which group’s treatment probability to increase, and which to decrease, when the objective is maximizing expected outcomes.
The next step, in Theorem 10, is to derive the population versions of the CADE and CAIE. These are purely population estimands, defined as the limit of $\bar{\tau}_{\text{CADE}}$ and $\bar{\tau}_{\text{CAIE}}$ as the market size grows to infinity. These estimands are a relevant policy counterfactual if we are interested in deploying a policy to a market that is much larger, but otherwise similar to the market being observed. The population estimands are also useful for deriving estimators that are consistent for both the large-market and finite-market estimands. Theorem 10 shows that the population estimands have a simple and interpretable form and implies that estimators for one class of estimands are consistent for the other.
Theorem 10.
Under the Assumptions of Theorem 2, and the additional assumption that for all $x\in\mathcal{S}$, $y(w,p,x)$ and $z(w,p,x)$ is continuously differentiable in $p$ for all $p\in\mathcal{S}$. Then, the population conditional average direct effect is:
$\lim\limits_{n\rightarrow\infty}\bar{\tau}_{\text{CADE}}(x)=\tau^{*}_{\text{CADE}}(x)=y(1,p^{*}_{\pi},x)-y(0,p^{*}_{\pi},x).$ (29)
The population conditional average indirect effect is:
$\tau^{*}_{\text{CAIE}}(x)=-\xi_{y}^{\top}\xi_{z}^{-1}\tau^{*,z}_{\text{CADE}}(x),\ \ \ \ \tau^{*,z}_{\text{CADE}}(x)=z(1,p^{*}_{\pi},x)-z(0,p^{*}_{\pi},x),$ (30)
and is the limit of $\bar{\tau}_{\text{CAIE}}(x)$, where the convergence is over all sets with positive measure:
$\lim\limits_{n\rightarrow\infty}\mathbb{E}\left[\bar{\tau}_{\text{CAIE}}(X)-\tau^{*}_{\text{CAIE}}(X)\,\big{|}\,X\in S\right]=0$ (31)
for all sets $S\subseteq\mathcal{X}$ such that $\mathbb{P}\left[X\in S\right]>0$.
The population estimands have a simple form that suggest estimation strategies for the estimands $\bar{\tau}_{\text{CADE}}(x)$ and $\bar{\tau}_{\text{CAIE}}(x)$. The limit of the CADE is the direct treatment effect on outcomes conditional on $x$, holding the equilibrium price fixed. The limit of the CAIE is the direct treatment effect on net demand conditional on $x$, multiplied by an elasticity correction that does not depend on $x$. The augmented randomized experiment from Design 2 can be used to estimate the elasticity corrections and the conditional average treatment effects required to estimate $\bar{\tau}_{\text{CADE}}(x)$ and $\bar{\tau}_{\text{CAIE}}(x)$.
For the elasticity corrections, the market price is an aggregate statistic, so individuals with different covariates all respond to the same market prices. A change in net demand of a given size always has the same impact on the market price. Although individuals’ responses to the treatment through outcomes or net demand are heterogeneous, the elasticity correction that transforms the $\tau_{\text{CADE}}^{*,z}(x)$ to $\tau_{\text{CAIE}}^{*}(x)$ is unconditional. This implies that a group of individuals’ effect on the system depends on their covariates only through conditional direct effects. Estimators for $\xi_{y}$ and $\xi_{z}$ from the previous section of the paper apply directly.
Below, we show that the $k$-nearest neighbor estimator is consistent for $\tau_{\text{CADE}}^{*}(x)$; the same result also immediately holds for $\tau_{\text{CADE}}^{*,z}(x)$. The salient fact in establishing this result is that the $k$-nearest neighbor estimator is a difference-in-means estimator that has been localized in $X$-space; and the proof suggests that other standard CATE estimators that are effectively localized difference-in-means estimators, such as the causal trees of Athey & Imbens (2016), are also consistent for the CADE. In our experiments, we use causal forests as implemented in the grf package of Athey et al. (2019) to estimate the CADE.
Theorem 11.
Let $N_{k}(x)$ be the $k$ closest observations to $x$ in terms of covariate distance $\left\lVert X_{i}-x\right\rVert_{2}$, breaking ties randomly if needed. Suppose that we collect data under Design 1, and construct the $k$-nearest neighbor estimator for $\tau_{\text{CADE}}(x)$ as
$\hat{\tau}_{\text{CADE}}(x)=\frac{\sum_{\left\{i\in N_{k}(x):W_{i}=1\right\}}Y_{i}}{\left\lvert\left\{i\in N_{k}(x):W_{i}=1\right\}\right\rvert}-\frac{\sum_{\left\{i\in N_{k}(x):W_{i}=0\right\}}Y_{i}}{\left\lvert\left\{i\in N_{k}(x):W_{i}=0\right\}\right\rvert}.$ (32)
Under the assumptions of Theorem 10, suppose furthermore that the conditional distribution of $\left\{Y_{i}(w,\,p),\,Z_{i}(w,\,p)\right\}$ given $X_{i}=x$ varies continuously in $x$. Then, given any sequence $k\rightarrow\infty,\,k/n\rightarrow 0$, the $k$-nearest neighbor estimator is consistent,
$\hat{\tau}_{\text{CADE}}(x)\rightarrow_{p}\tau_{\text{CADE}}^{*}(x),$ (33)
at any point $x\in\mathcal{X}$ with positive local mass, i.e., with $\mathbb{P}\left[\left\lVert X_{i}-x\right\rVert_{2}\leq\varepsilon\right]>0$ for any $\varepsilon>0$.
4.2 Equilibrium-Stable Targeting
The results of the previous section imply that estimates of $\tau_{\text{CADE}}^{*}(x)$ and $\tau_{\text{CAIE}}^{*}(x)$ could be used to (locally) optimize an unconstrained targeting policy. There are a variety of constrained targeting rules that may be of interest. Here, we focus on one specific question of this type, namely what is the optimal treatment assignment policy that does not move equilibrium prices relative to those seen in the experiment? There are two reasons to consider this class of targeting rules. First, a conservative policymaker may be reluctant to significantly modify the equilibrium—even if it is beneficial on average to individuals—and so they may want to know how much they can improve outcomes without changing the equilibrium. Second, from a practical point of view, we find that the answer to this question admits a simple econometric strategy, and can be answered without needing to estimate price elasticities and without recourse to an augmented experimental design. Instead, the optimal equilibrium-stable targeting rule (and its performance) can be estimated using data from a baseline RCT following Design 1, as long as the RCT collects outcome and relevant supply and demand data at an individual level.
The optimization problem for the equilibrium-stable policy is to find a new policy $\nu(\cdot)$ that maximizes expected outcomes in the population, while maintaining the population equilibrium price obtained under the policy $\pi(\cdot)$, i.e., $p^{*}_{\pi}$. The value of a given policy $\nu(\cdot)$ is:
$V(\nu)=\lim_{n\rightarrow\infty}\mathbb{E}\left[\mathbb{E}_{\nu}\left[Y_{i}\right]\right]=\mathbb{E}\left[(1-\nu(X_{i}))Y_{i}(0,\,p^{*}_{\nu})+\nu(X_{i})Y_{i}(1,\,p^{*}_{\nu})\right],$ (34)
where $\mathbb{E}_{\nu}\left[Y_{i}\right]$ denotes expected rewards under the considered new policy $\nu(\cdot)$. The optimal policy is defined as
[…]
where $p^{*}_{\nu}$ is the population equilibrium price under $\nu(\cdot)$.
When the targeting policy is restricted to have the same equilibrium effect as a baseline policy, then solving for that optimal policy takes the form of a linear optimization problem.
Proposition 12.
Under the assumptions of Theorem 2, optimizing the target (34) across all asymptotically equilibrium-stable policies is equivalent to solving the following linear optimization problem:
$\nu^{*}(\cdot)=\stackrel{{\scriptstyle[}}{{\nu\,:\,\mathcal{X}\rightarrow[0,\,1]}}]{}{\operatorname{argmax}}\left\{\mathbb{E}\left[\nu(X_{i})\,\tau^{*}_{\text{CADE}}(X_{i})\right]:\mathbb{E}\left[\left(\nu(X_{i})-\pi(X_{i})\right)\tau^{*,z}_{\text{CADE}}(X_{i})\right]=0\right\}.$ (36)
The above reveals that we can solve for the optimal price-stable policy without access to price elasticities. Thus, we can obtain a plug-in estimate for the optimal rule using CADE estimates derived from an RCT without price perturbations as in Design 1; recall that the CADE itself can be identified without price perturbations (Theorem 11).
The characterization in Proposition 12 involves a functional linear program. We show below, however, that its solution takes on a simple parametric form: The optimal price-stable policy is a thresholding rule that compares the CADE for the outcomes against a shadow-cost of net demand effects $c\cdot\tau_{\text{CADE}}^{*,z}(x)$, where $c\in\mathbb{R}^{J}$ can be interpreted as a shadow price vector.
Theorem 13.
Under the conditions of Proposition 12, the asymptotic equilibrium-stable targeting problem (36) admits an optimal solution of the form
$\nu_{c,a}(x)=\begin{cases}1&\text{if}\ \ \tau^{*}_{\text{CADE}}(x)>c^{\top}\,\tau_{\text{CADE}}^{*,z}(x),\\ a&\text{if}\ \ \tau^{*}_{\text{CADE}}(x)=c^{\top}\,\tau_{\text{CADE}}^{*,z}(x),\\ 0&\text{else,}\end{cases}$ (37)
for some $c\in\mathbb{R^{J}}$ and $a\in[0,1]$.
The targeting rule in Theorem 13 takes on a particularly simple form when we only have $J=1$ good, and the intervention has a crowding-out type structure where direct effects are positive but spillovers are negative, $\tau^{*}_{\text{CADE}}(x)>0$ and $\tau_{\text{CADE}}^{*,z}(x)<0$. In this case, the optimal price-stable targeting rule will be of the form
$\nu^{*}(x)=1\left(\left\{\frac{\tau^{*}_{\text{CADE}}(x)}{-\tau^{*,z}_{\text{CADE}}(x)}>c\right\}\right)$ (38)
with potential random tie-breaking at the cutoff. In other words, we want to give treatment to units whose direct effects are large relative to corresponding spillovers; we then pick $c$ to satisfy the price-stability constraint. This algorithmic structure is familiar from the literature on cost-sensitive treatment targeting (e.g., Sun et al. 2021).
5 Example: Cash-Transfer Experiments
In this section, we estimate a simple model of the direct and indirect impacts of a conditional cash transfer on children’s health outcomes in the Philippines using data from Filmer et al. (2023). We use data simulated from the model to illustrate the performance of the average direct and indirect treatment effect estimators and the estimated targeting rule proposed in this paper. Filmer et al. (2023) use a village-level, cluster-randomized experiment to analyze the effect of the Pantawid conditional cash transfer program on young children’s health outcomes. The program provides a monthly sum for eligible families, where the amount depends on the number of children in the household. Families are eligible if they meet a proxy means test and have either children under 14 years of age or a pregnant woman in the household. The authors show that the health of eligible children in treated villages is better than in control villages. However, children whose families are ineligible for the transfer are worse off in treated villages compared to control villages.
The authors argue that this is phenomenon is due to spillover effects, which largely arise from equilibrium effects in the market for perishable protein sources. Specifically, the authors provide evidence that eggs, which are a commonly consumed perishable protein source in the Philippines, have higher prices in treated compared to control villages where saturation of eligible households is high. This price effect is explained by the increase in demand for protein from eligible families and inelastic supply in remote villages. The higher prices in treated villages lead ineligible children to consume less protein, which explains a significant proportion of the negative effect of the cash transfer program on ineligible children. It thus appears that Assumption 1 is reasonable, in that the spillovers described in detail by the authors are restricted to those that occur through market prices.
5.1 Treatment Effects in a Village Economy
Under the assumption that markets in remote villages are isolated from one another, the village-level randomized experiment allowed Filmer et al. (2023) to recover treatment effects that include equilibrium effects. The total effect $\tau_{\text{TOT}}$ could be estimated using data from the cluster-randomized experiment by comparing the average outcomes of all children in treated compared to control villages. In this paper, in contrast, our focus is on randomized experiments where only a single market is observed, but randomization is possible at the household level. While our econometric setting is different from that of Filmer et al. (2023), we can use their dataset to conduct a calibrated evaluation of our approach. We first use moments from the cluster-randomized experiment Filmer et al. (2023) to estimate a model of supply, demand, and children’s outcomes in a single remote village market. Then, we use this model to simulate a single-village, unit-randomized experiment, and assess the ability of our estimators of the average direct and indirect treatment effects to recover equilibrium effects from this simulated experiment.
Let $i\in\{1,\,\ldots,n\}$ index the people in the village and $h\in\{1,\,\ldots n_{h}\}$ index the households in the village. For each $h\in\{1,\,\ldots,n_{h}\}$, $A^{h}$ is a set of integers containing the indexes of the people that are part of household $h$ and $C^{h}$ is a set of integers containing the indexes of the children aged 0-5 that are part of household $h$. $h(\cdot)$ is a map that provides the household index given a person’s index. The binary treatment $W_{h}\in\{0,1\}$ is assigned at the household level. $E_{h}\in\{0,1\}$ indexes the eligibility for a household. We assume that individual demand in eggs per week is determined by the following linear equation, where $p$ is the price of eggs:
$D_{i}(W_{h},p)=\theta_{d01}\cdot E_{h(i)}+\theta_{d00}\cdot(1-E_{h(i)})+\theta_{dw}\cdot W_{h(i)}\cdot E_{h(i)}+\theta_{dp}\cdot p+\varepsilon_{d,h(i)}+\nu_{d,i}(W_{i}),$
where the household and individual-level error terms are normally distributed. Demand for eggs is increased for eligible adults and children in households that receive the Pantawid conditional cash transfer. Demand also responds to the market price of eggs.
We assume that the aggregate supply per person in the village is responds linearly to prices, $S(p)=\theta_{s0}+\theta_{sp}\cdot p$. This formulation assumes that the village supply is not directly affected by the treatment; in Filmer et al. (2023), the authors argue that egg supply in the Philippines is dominated by larger commercial producers, which makes this a reasonable assumption.
Our target outcome is the height-for-age Z-score for children between 0 and 5 years of age. For our evaluation, we assume that these outcomes are generated as
$Y_{i}(W_{h(i)},p)=\theta_{y01}E_{h(i)}+\theta_{y00}(1-E_{h(i)})+\theta_{yd}D_{i}(W_{h(i)},p)+\theta_{yw}W_{h(i)}+\varepsilon_{y,h(i)}+\nu_{yi}(W_{i}),$
where the household and individual-level error terms are normally distributed. Consumption of a perishable protein source increases child-level health outcomes. The conditional cash transfer also has impact on health metrics through other factors, which is captured by $\theta_{yw}$.
For the purpose of our analysis, we aggregate individual level outcomes at the household level:
$D_{h}(W_{h},p)=\frac{n_{h}}{n}\sum\limits_{i\in A^{h}}D_{i}(W_{h},p),\qquad Y_{h}(W_{h},p)=\frac{n_{h}}{n_{c}}\sum\limits_{i\in C^{h}}Y_{i}(W_{h(i)},p).$ (39)
The factor $\frac{n_{h}}{n}$ ensures that average treatment effects on household-level demand can be interpreted as the treatment effect on eggs consumed per-person in the village. The factor $\frac{n_{h}}{n_{c}}$ ensures that average treatment effects on household-level outcomes can be interpreted as the treatment effect per child aged 0 to 5 in the village. The realized egg price in the market matches aggregate supply and demand:
$\frac{1}{n_{h}}\sum\limits_{h=1}^{n_{h}}D_{h}(W_{h(i)},P(\bm{W}))=S(P(\bm{W}))$
Given a treatment vector $\bm{W}$, the realized household-level outcomes are $Y_{h}(\bm{W})=Y_{h}(W_{h},P(\bm{W}))$ and household level net demand is $Z_{h}(\bm{W})=D_{h}(\bm{W},P(\bm{W}))-\frac{1}{n_{h}}S(P(\bm{W}))$.
The 10 parameters in this model are identified by 10 moments from a subset of the data in Filmer et al. (2023), restricted to households in villages above the 70th percentile of the remoteness index defined in the paper, where equilibrium effects were most meaningful. The estimated model matches these 10 moments:
  • Average outcomes and demand for eggs for eligible and ineligible children in control villages;
  • The average price of eggs in treated villages and control villages;
  • The elasticity of demand for eggs in the Philippines reported in the paper;
  • The total treatment effect on outcomes and demand for eligible children;
  • The ratio of total treatment effects on outcomes and demand for ineligible children.
The total treatment effect here is the difference in average outcomes (or demand) in treated versus control villages. The estimation details and the estimated parameters of the model are reported in Appendix A.
Figure 1(a) shows that aggregate demand in the village is decreasing in price, but is shifted upward by the cash transfer when eligible families are treated. Supply is increasing in price but is inelastic enough to explain the shift in prices observed in the data in treated villages. Figure 1(b) shows the distribution of $\tau_{\text{ADE}}$ and $\tau_{\text{TOT}}$, simulated using repeated samples of a village of 1000 households from the model. In this linear model, $\tau_{\text{TOT}}=\tau_{\text{MPE}}$. An RCT estimates the direct effect, which is more than 50% higher than the total effect. Figure 1(a) explains why the direct effect and the total effect are so different in this setting. The treatment raises demand for eggs among treated families, which impacts the market-clearing price for the treated market compared to the control market. Holding food prices fixed, the effect of the treatment is a large positive increase in health outcomes. However, when the effect on food prices is taken into account, which impacts both eligible and ineligible children, the total treatment effect is much lower.
(a) Estimated expected demand and supply under treatment and control. Prices are per egg in local currency.
(b) Smoothed empirical distribution of the average direct effect $\tau_{\text{ADE}}$ and the total effect $\tau_{\text{TOT}}$ for samples of a village of $1,000$ households.
Figure 1: Illustration of the gap between $\tau_{\text{ADE}}$ and $\tau_{\text{TOT}}$ in the village market model. The intervention has both a direct effect on children’s health and an offsetting indirect effect through increased food prices.
In Table 1, we show that both the direct and the indirect effect are estimable in finite samples using our proposed unit level augmented experiment. We report the results of a Monte Carlo simulation with 1,000 repetitions and a sample size of 2,000 households to evaluate the bias, variance, and coverage properties of the estimators and confidence intervals for $\hat{\tau}_{\text{ADE}}$ and $\hat{\tau}_{\text{AIE}}$. The table illustrates that the bias of the average direct effect and indirect effect estimator based on the augmented randomize experiment are low in finite samples. Furthermore, the coverage when $\hat{\sigma}_{D}$ and $\hat{\sigma}_{I}$ are used to construct confidence intervals in finite samples is slightly above the asymptotic confidence level of 95% for both population estimands. As expected from Theorem 3, intervals that target $\tau^{*}_{\text{ADE}}$ are conservative for $\tau_{\text{ADE}}$. For the AIE, in simulations the coverage for the sample estimand is also more conservative than for the population estimand.
Estimate Bias S.D. Coverage for $\tau$ Coverage for $\tau^{*}$
$\hat{\tau}_{\text{ADE}}$ 0.317 -0.005 0.123 0.965 0.959
$\hat{\tau}_{\text{AIE}}$0.132 0.002 0.082 0.964 0.961
Table 1: Monte Carlo Simulation Results for $\hat{\tau}_{\text{ADE}}$ and $\hat{\tau}_{\text{AIE}}$ for a sample of $n=2,000$ and $1,000$ repetitions
5.2 Heterogeneous Treatment Effects in the Model
In the setting of Filmer et al. (2023), there may be heterogeneity in both outcome and demand effects that is correlated with observed pre-treatment covariates. If so, a targeting rule can improve outcomes compared to a randomized rule that induces the same equilibrium. Since the replication package did not contain data on pre-treatment household characteristics that can be matched to post-treatment data, we cannot estimate heterogeneous treatment effects directly. Instead, we augment the model in Section 5.1 with an additional 10 generated household-level covariates that are observed by the planner, and a set of four heterogeneous household types that are correlated with the generated covariates but unobserved. These household types are generated purely for this simulation exercise and are not calibrated to real data. Household Type A has the same data-generating process in the previous section. Household Type B is especially focused on their children, and purchases enough food that outcomes improve even more than for Type A. For Household Type C, which makes up a small percentage of the households, increased cash is not beneficial for child health outcomes, as it makes consuming some welfare-decreasing product possible for adults, and the budget share of food of the household decreases. For Household Type D, there is a small positive effect on children’s health outcomes, but a large effect on demand for the perishable protein source; in these houses, adults consume most of the additional food that is purchased with the cash transfer. The details of the augmented model are in Appendix A. Then, we estimate the optimal equilibrium-stable targeting rule on data simulated from this model.
In Figure 2, we illustrate the optimal rule estimated using a sample of $5,000$ households. Ignoring equilibrium effects, a good targeting rule would allocate treatment to those individuals who are estimated to have a positive CADE. Here, however, this rule would result in a larger impact on perishable food demand compared to the RCT since, on average, those who increase their children’s health using the cash transfer also increase their consumption of perishable food. In contrast, the optimal rule that respects the equilibrium constraint allocates treatment to those with $\hat{\tau}_{\text{CADE}}(X_{i})>c\hat{\tau}^{z}_{\text{CADE}}(X_{i})$. We can see that for the most part, it is those with a positive CADE that are treated. However, households of Type D consume a lot of additional perishable food without raising a child’s health outcomes very much. Dropping some of these households from the equilibrium-stable targeting rule allows the planner to target more households from Type B while keeping the equilibrium food price stable.
Figure 2: A scatter-plot of an estimate of $\tau_{\text{CADE}}(X_{i})$ and $\tau^{z}_{\text{CADE}}(X_{i})$ for a sample of 125 households, estimated using the causal forest of Athey et al. (2019). Overlaid is the optimal equilibrium-stable targeting rule, computed by solving a plug-in version of the linear program in Proposition 12 on a sample of 5,000 households, which takes the form of the rule in Theorem 13.
Optimal Rule Direct Rule Randomized Rule
Estimated Height-For-Age Z-Score -1.749 -1.925 -2.112
Standard Deviation 0.09 0.092 0.026
Table 2: Value of estimated equilibrium-stable targeting rule. Reported values are average and standard deviation across 50 samples of 5,000 households drawn from the model.
We then evaluate the optimal equilibrium stable rule by estimating the rule using the procedure in Section 4.2 on a training sample of 5,000 households, and then evaluating the expected value of that rule numerically on a separated simulated dataset of 5,000 households. The results are reported in Table 2. We compare the outcomes under this rule to two baselines. The first baseline is the randomized experiment, where 55% of eligible households are treated. The second is an equilibrium-stable version of a treatment rule that targets based on $\hat{\tau}_{\text{CADE}}(x)$ only. This rule is estimated on the training sample as $\tilde{\pi}(x)=\hat{\alpha}\cdot\mathbbm{1}(\hat{\tau}_{\text{CADE}}(x)>0)$. $\hat{\alpha}$ is chosen so that, in the training data, the total net demand under $\tilde{\pi}$ is equal to the total net demand in the randomized experiment. This rule avoids treating some of the few households for which the cash transfer is not beneficial, which improves outcomes compared to a purely randomized rule. However, for equilibrium stability, $\alpha<1$, and the rule does not distinguish between households based on their demand for the perishable protein source. The optimal rule has higher outcomes since it treats more households with a large positive direct effect on outcomes and avoids treating households for which the treatment has a small impact on outcomes but a large impact on demand.
6 Discussion
Analyzing the performance of randomized control trials in settings with equilibrium effects is needed given the rapid growth of experimentation both in practice and in research studies. The Neyman-Rubin framework that relies on SUTVA rules out interaction effects that can have an important impact on decision-relevant treatment effects. A parametric structural model may capture a variety of complex equilibrium effects, but is not robust to misspecification, which can be problematic when individuals behave in complex and heterogeneous ways. A model of treatment effects under general patterns of interference is intractable without clustering or other assumptions.
This paper shows that it is fruitful to marry ex-ante knowledge about the structure of an economic environment with a non-parametric stochastic model of treatment effects under interference. This leads to a characterization of asymptotic properties of treatment effects and estimators of those treatment effects based on new and existing experimental designs that are robust to a wide range of modeling choices. Results on estimation, inference, and optimal targeting in complex environments with some economic structure imposed can then be easily compared with the large body of work that studies causal inference under SUTVA.
Our results on the direct effect and on targeting rely on the assumption that all spillovers are mediated by the prices of a finite number of traded goods; our results for the indirect effect and for inference further rely on us being able to exogenously apply small shifts to the prices that different market participants are exposed to. As discussed in the previous section, such assumptions may be relevant to experiments run in communities with self-contained markets. Aouad et al. (2024) describe an experimental setting where researchers open a subsidy store that they then use to randomize subsidy eligibility at the unit level; and experiments of this type could plausibly also integrate small, random price perturbations. Another class of applications our methods are a natural fit for are online platforms for ride sharing, freelance labor, short-term rentals, etc. Price-mediated spillovers are of central importance when experimenting in such systems; for example, in ride sharing, both drivers and riders respond to the average price of a mile and/or minute of transportation, and both supply- and demand-side interventions may alter the market-clearing price for transportation. Furthermore, technology companies running such platforms have already documented willingness to run experiments comparable to the price perturbations we describe, e.g., via random discounts or bonuses (Castillo 2023, Holtz et al. 2024). There remains, however, a large class of settings where our methods cannot be applied, e.g., when spillovers are not mediated via observed equilibrium quantities (e.g., Cai et al. 2015, Manski 1993), or in matching problems where equilibrium dynamics play a key role (Li et al. 2023).
There are a variety of avenues for future work possible. One limitation of our approach is that we analyze the large sample limit of the market place where the number of suppliers and the number of buyers grow large; an analysis of experiments in settings where firms have significant market power would likely require different techniques. A related potential direction for extending our results would be to allow for the market to depend on prices of a continuum of goods, e.g., localized prices that vary continuously in space. In general, extending our results to a broader class of equilibrium mechanisms would be of considerable interest.
Appendix A Details from the Simulations in Section
A.1 Estimation
The model described in the main text captures the impact of a conditional cash transfer on children’s health outcomes and the market for a perishable protein source in a remote village in the Philippines. It is designed based on the results in Filmer et al. (2023). The household and individual-level error terms in the demand equation have standard deviation of 1/3rd. The household level error terms in the outcome equation have standard deviation of 1/3rd and the individual-level error terms have standard deviation of 1.
There are 10 coefficients of the model $\bm{\theta}=\begin{bmatrix}\theta_{d00}&\theta_{d01}&\theta_{dw}&\theta_{dp}&\theta_{s0}&\theta_{sp}&\theta_{y00}&\theta_{y01}&\theta_{yd}&\theta_{yw}\end{bmatrix}^{\top}$. The model is estimated to match the following 10 moments from the subsample of remote villages in Filmer et al. (2023). These moments are computed using data files found in the replication package for the paper, see Filmer et al. (2021).
  • The average demand for eggs per week among eligible and ineligible children in control villages $(\hat{\mu}^{d}_{elig},\hat{\mu}^{d}_{inelig})$.
  • The average height-for-age Z-score among eligible and ineligible children in control villages ($\hat{\mu}^{y}_{elig},\hat{\mu}^{y}_{inelig})$.
  • The average price of eggs in treated and control villages ($\hat{p}^{0}$ and $\hat{p}^{1}$). The elasticity of demand for individuals in control villages in the model matches the elasticity of egg demand cited, although not directly estimated in Filmer et al. (2023) $(\eta_{egg})$.
  • The effect of the treatment on eligible children on height-for-age Z-score and eggs per week ($\hat{\tau}^{y}_{elig},\hat{\tau}^{d}_{elig})$, where treatment effects are defined as the differences-in-means for ineligible children in treated versus control villages.
  • The ratio of treatment effects on height-for-age Z-score and eggs-per-week demand for ineligible children ($\hat{\pi}=\hat{\tau}^{y}_{inelig}/\hat{\tau}^{d}_{inelig}$) [^6] It is possible to estimate the model by matching the treatment effects on outcomes and demand directly for ineligible children, rather than their ratio. This would avoid using the elasticity of egg demand, which is taken from the literature rather than estimated from the data. However, the sample size of ineligible children in remote villages is small, so estimating the model to match these individual moments, which are imprecisely estimated, rather than their ratio, leads to an implied elasticity that does not seem reasonable.. Again, treatment effects are defined as the differences-in-means for ineligible children in treated versus control villages.
The model is just-identified by using these moments. Each parameter has an estimator, which is a simple function of these moments:
$\begin{split}&\hat{\theta}_{d00}=\hat{\mu}^{d}_{inelig}-\hat{\theta}_{dp}\cdot\hat{p}^{0},\qquad\hat{\theta}_{d01}=\hat{\mu}^{d}_{elig}-\hat{\theta}_{dp}\cdot\hat{p}^{0},\\ &\hat{\theta}_{dp}=\frac{\eta_{egg}}{1.01\cdot\hat{p}_{0}}\cdot\frac{1}{n}\sum\limits_{i=1}^{n}(E_{i}\hat{\mu}^{d}_{elig}+(1-E_{i})\hat{\mu}^{d}_{inelig}),\qquad\hat{\theta}_{dw}=\hat{\tau}^{d}_{elig}-\hat{\theta}_{dp}(\hat{p}_{1}-\hat{p}_{0}),\\ &\hat{\theta}_{sp}=\left(\frac{1}{n}\sum\limits_{i=1}^{n}E_{i}\hat{\tau}^{d}_{elig}+\frac{1}{n}\sum\limits_{i=1}^{n}(1-E_{i})(\hat{p}^{1}-\hat{p}^{0})\hat{\theta}_{dp}\right)\Big{/}(\hat{p}^{1}-\hat{p}^{0})\\ &\hat{\theta}_{s0}=\frac{1}{n}\sum\limits_{i=1}^{n}(E_{i}\hat{\mu}^{d}_{elig}+(1-E_{i})\hat{\mu}^{d}_{inelig})-\hat{\theta}_{sp}\hat{p}_{0},\\ &\hat{\theta}_{yd}=\frac{\hat{\tau}^{y}_{inelig}}{\hat{\tau}^{d}_{inelig}},\qquad\hat{\theta}_{yw}=\hat{\tau}^{y}_{elig}-\hat{\theta}_{yd}\hat{\tau}^{d}_{inelig},\\ &\hat{\theta}_{y00}=\hat{\mu}^{y}_{inelig}-\hat{\theta}_{yd}\hat{\mu}^{d}_{inelig},\qquad\hat{\theta}_{y01}=\hat{\mu}^{y}_{elig}-\hat{\theta}_{yd}\hat{\mu}^{d}_{inelig}.\end{split}$
The moments used for estimating the model and the corresponding parameter estimates are described in Table 3.
Estimated Moments from Filmer et al. (2023)
$\hat{\mu}^{d}_{elig}$ $\hat{\mu}^{d}_{inelig}$ $\hat{\mu}^{y}_{elig}$ $\hat{\mu}^{y}_{inelig}$ $\hat{p}^{0}$ $\hat{p}^{1}$ $\eta_{egg}$ $\hat{\tau}^{y}_{elig}$ $\hat{\tau}^{d}_{elig}$ $\frac{\hat{\tau}^{y}_{inelig}}{\hat{\tau}^{d}_{inelig}}$
1.6 2.2 -2.47 -1.44 6.07 6.26 -1.3 0.32 0.12 1.85
Estimated Parameters
$\hat{\theta}_{d00}$ $\hat{\theta}_{d01}$ $\hat{\theta}_{dw}$ $\hat{\theta}_{dp}$ $\hat{\theta}_{s0}$ $\hat{\theta}_{sp}$ $\hat{\theta}_{y00}$ $\hat{\theta}_{y01}$ $\hat{\theta}_{yd}$ $\hat{\theta}_{yw}$
4.49 3.89 0.19 -0.38 -0.21 0.33 -5.51 -5.43 1.85 0.1
Table 3: Estimation of the Model in Section 3
Last, in drawing data from the model, $C_{i}$ and $h(i)$ are determined by random draws from the empirical distribution of households with children between 0 and 14 (who are in eligible for the transfer if they meet a proxy means test) from the baseline survey in Filmer et al. (2023). $E_{i}$ is drawn randomly to match the average percentage of eligible households in remote villages.
A.2 Adding Heterogeneity
The model in Section 5.1 is augmented to add some additional heterogeneity. For each household, a set of covariates $X_{h}\sim\mathcal{N}(0,1)^{10}$ is drawn. Households now have type $\zeta_{h}\in\{A,B,C,D\}$. Households of type $\zeta_{h}=A$ have data generated by the model in the previous section. In response to the treatment, they increase their purchase of a perishable protein source, and children’s health outcomes are improved on average.
$\begin{split}&D^{A}_{i}(W_{h},p)=\theta_{d01}\cdot E_{h(i)}+\theta_{d00}(1-E_{h(i)})+\theta_{dw}W_{h(i)}E_{h(i)}+\theta_{dp}\cdot p+\varepsilon_{d,h(i)}+\nu_{d,i}(W_{i})\\ &Y^{A}_{i}(W_{h(i)},p)=\theta_{y01}E_{h(i)}+\theta_{y00}(1-E_{h(i)})+\theta_{yd}D^{A}_{i}(W_{h(i)},p)+\theta_{yw}W_{h(i)}E_{h(i)}+\varepsilon_{y,h(i)}+\nu_{yi}(W_{i})\end{split}$
For the other three types, treatment effects are distinct from Type A. Households of type $\zeta_{h}=B$ increase their consumption of a perishable protein source and purchase a more expensive non-perishable protein source (such as canned fish). We assume for this second type of good, because it is non-perishable, supply is much more elastic and equilibrium effects are negligible. So, for these households, the data generating process is as follows:
$\begin{split}&D^{B}_{i}(W_{h},p)=D^{A}_{i}(W_{h},p)+0.75\theta_{d01}\cdot E_{h(i)}\cdot W_{h(i)}\\ &Y^{B}_{i}(W_{h(i)},p)=\theta_{y01}E_{h(i)}+\theta_{y00}(1-E_{h(i)})+\theta_{yd}D^{A}_{i}(W_{h(i)},p)+(\theta_{yw}+0.2)W_{h(i)}E_{h(i)}+\varepsilon_{y,h(i)}+\nu_{yi}(W_{i})\end{split}$
For households of type $\zeta_{h}=C$, the cash transfer does not have its intended positive effect. In response to the cash transfer, an increased ability of parents to afford addictive goods (such as cigarettes, or alcohol) leads to a decrease in the purchase of perishable protein, which has a large negative effect on children’s health. Let $\mbox{Child}_{i}$ be an indicator if individual $i$ is under 5 years of age.
$\begin{split}&D^{C}_{i}(W_{h},p)=\theta_{d01}\cdot E_{h(i)}(1-0.75W_{h(i)}\mbox{Child}_{i})+\theta_{d00}(1-E_{h(i)})+\theta_{dp}\cdot p+\varepsilon_{d,h(i)}+\nu_{d,i}(W_{i})\\ &Y^{C}_{i}(W_{h(i)},p)=\theta_{y01}E_{h(i)}+\theta_{y00}(1-E_{h(i)})+\theta_{yd}D_{i}(W_{h(i)},p)+\theta_{yw}W_{h(i)}E_{h(i)}+\varepsilon_{y,h(i)}+\nu_{yi}(W_{i})\end{split}$
For households of type $\zeta_{h}=D$, there is an increase in food consumption of adults, but not of children.
$\begin{split}&D^{D}_{i}(W_{h},p)=\theta_{d01}\cdot E_{h(i)}+\theta_{d00}(1-E_{h(i)})+\theta_{dp}\cdot p+\varepsilon_{d,h(i)}+1.5\theta_{dw}W_{h(i)}E_{h(i)}(1-\mbox{Child}_{i})+\nu_{d,i}(W_{i})\\ &Y^{D}_{i}(W_{h(i)},p)=\theta_{y01}E_{h(i)}+\theta_{y00}(1-E_{h(i)})+\theta_{yd}D_{i}(W_{h(i)},p)+3\cdot\theta_{yw}W_{h(i)}E_{h(i)}+\varepsilon_{y,h(i)}+\nu_{yi}(W_{i})\end{split}$
The household type is correlated with observed covariates as follows. The score for each household type is:
$U_{hA}$ $=1+X_{2h}-0.5X_{3h}+\varepsilon_{hA},$
$U_{hB}$ $=X_{1h}+\varepsilon_{hB},$
$U_{hC}$ $=-1+X_{3h}+\varepsilon_{hC},$
$U_{hD}$ $=X_{4h}+\varepsilon_{hD},$
where each of the noise terms are drawn from a Type I extreme value distribution. $\zeta_{h}=\arg\max\limits_{T\in\{A,B,C,D\}}U_{hT}$. The coefficients of the model are the same as in Table 3. In the simulation, 44% of the households are Type A, 22% are type B, 22% are Type D, and 12% are Type C. Demand and outcomes are aggregated by household as in Equation (39).
Appendix B Proof of Main Results
We use the following notation through the proof. For a deterministic sequence $X_{n}$ and a positive, deterministic sequence $a_{n}$, we say that $X_{n}=\mathcal{O}(a_{n})$ if $\limsup_{n\rightarrow\infty}X_{n}/a_{n}<\infty$, and $X_{n}=o(a_{n})$ if $\limsup_{n\rightarrow\infty}X_{n}/a_{n}<0$. For a sequence of random variables $X_{n}$, and $a_{n}$ as before, we say that $X_{n}=\mathcal{O}_{p}(\alpha_{n})$ if for every $\varepsilon$, there exists an $c<\infty$ such that $\lim\limits_{n\rightarrow\infty}\mathbb{P}\left[\|A_{n}\|>\alpha_{n}c\right]<\varepsilon$, and we say that $X_{n}=o(\alpha_{n})$ if, for any $c>0$, $\lim\limits_{n\rightarrow\infty}\mathbb{P}\left[\|X_{n}\|>\alpha_{n}c\right]=0$.
B.1 Proof of Theorem 1
We start by proving the following more general central limit theorem for market prices $P_{n}(\bm{W})$ in a system with potential price perturbations. Theorem 1 follows from Lemma 14.
Lemma 14.
Under the setting of Theorem 1, suppose furthermore that the market may have random price perturbations $U_{i}$ as in Design 2 with $\mathbb{E}\left[U_{i}\right]=0$ and $\left\lVert U_{i}\right\rVert\leq h_{n}$ almost surely with $h_{n}=o(n^{-1/4})$. Then,
$P_{n}(\bm{W})-p^{*}_{\pi}=-\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi})+o_{p}(n^{-1/2}).$ (40)
and furthermore $\sqrt{n}\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\Rightarrow\mathcal{N}\left(0,\,\operatorname{Var}\left[Z_{i}(W_{i},p^{*}_{\pi})\right]\right)$.
Proof.
Because $\mathcal{S}$ is compact, Assumptions 2 and 5 together imply that not only is there is a unique population equilibrium price $p^{*}_{\pi}$ with $z(p^{*}_{\pi})$, but there is also a critical radius $\varepsilon^{*}$ such that
$\left\lVert z(p)\right\rVert_{2}\geq\frac{\lambda_{\min}\left(\xi_{z}\right)}{2}\min\left\{\left\lVert p-p^{*}_{\pi}\right\rVert_{2},\,\varepsilon^{*}\right\}.$ (41)
To verify this result, we note that, by twice differentiability of $z$, there exists a $\delta^{*}>0$ such that
$\left\lVert z(p)\right\rVert_{2}\geq\frac{\lambda_{\min}\left(\xi_{z}\right)}{2}\left\lVert p-p^{*}_{\pi}\right\rVert_{2}\text{ for all }\left\lVert p-p^{*}_{\pi}\right\rVert_{2}\leq\delta^{*}.$
Thus, (41) must hold with
$\varepsilon^{*}=\min\left\{\delta^{*},\,\frac{2}{\lambda_{\min}\left(\xi_{z}\right)}\inf\left\{\left\lVert z(p)\right\rVert_{2}:\left\lVert p-p^{*}_{\pi}\right\rVert_{2}>\delta,\,p\in\mathcal{S}\right\}\right\},$
where the latter term must because $\mathcal{S}$ is compact, $z(p)$ is continuous over $\mathcal{S}$ and $p^{*}_{\pi}$ is the unique equilibrium point over $\mathcal{S}$.
We next use this fact to establish tail bounds on $\sqrt{n}\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert_{2}$; we will then sharpen these into a central limit theorem in a second step. Write
[…]
By Lemma 19, there is a constant $C$ such that, for all $t\geq 1$,
[…]
Furthermore, because $z(p)$ is twice continuously differentiable in $p$ and we have $\mathbb{E}\left[U_{i}\right]=0$ and $\left\lVert U\right\rVert\leq h_{n}$ almost surely
[…]
Thus, for $n$ large enough,
[…]
$\mathbb{P}\left[\left\lVert z(P_{n}(\bm{W}))\right\rVert>\frac{2+t}{\sqrt{n}}\right]\leq Ct^{2J}e^{-2t^{2}}+e^{-cn}.$
Together with (41), we then conclude that, for $n$ large enough, we have
$\mathbb{P}\left[\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert>\frac{1}{\sqrt{n}}\frac{4+2t}{\lambda_{\min}\left(\xi_{z}\right)}\right]\leq Ct^{2J}e^{-2t^{2}}+e^{-cn}$ (43)
for all $t\leq\sqrt{n}\lambda_{\min}\left(\xi_{z}\right)\varepsilon^{*}/2-2$, which in particular implies that $P_{n}(\bm{W})=p^{*}_{\pi}+\mathcal{O}_{p}(1/\sqrt{n})$.
Now, to sharpen our result into a central limit theorem, we invoke Lemma 20 to verify that, because $P_{n}(\bm{W})$ is consistent for $p^{*}_{\pi}$, we have asymptotic equicontinuity
[…]
Together with (42) this becomes
[…]
Again invoking twice continuous differentiability of $z(p)$, we get
[…]
Finally recall that, by Assumption 3, $\operatorname{Var}\left[Z_{i}(W_{i},\,p^{*}_{\pi})\right]\succ 0$ and so weak continuity implies that $\lim_{n\rightarrow\infty}\text{Corr}\left[Z_{i}(W_{i},\,p^{*}_{\pi}),\,Z_{i}(W_{i},\,p^{*}_{\pi}+U_{i})\right]=1$. The claimed central limit theorem then follows by the standard Lindeberg argument thanks to boundedness of $Z_{i}(w,\,p)$. ∎
B.2 Proof of Theorems 2, 3 and 4
We structure the bulk of the proof in terms of two lemmas given below, which also allow for potential price perturbations in anticipation of later needs. Given these two results, we first note that Theorem 3 is in fact a special case of Lemma 15 with $U_{i}=0$ almost surely. Moreover, the claim about the ADE in Theorem 2 follows immediately from (11) by noting that $\tau_{\text{ADE}}$ is almost surely bounded (because the outcomes are), and that $\mathbb{E}\left[Z_{i}(W_{i},\,p^{*}_{\pi})\right]=0$.
For our results about the AIE, we rely on a result of Hu et al. (2022, Theorem 1), who show that in Bernoulli trials with independent $W_{i}$ (but not necessarily constant treatment probabilities $\pi_{i}$), we have
$\tau_{\text{MPE}}=\tau_{\text{ADE}}+\tau_{\text{AIE}}.$ (47)
Lemmas 16 shows that $\tau_{\text{MPE}}$ converges to $\tau^{*}_{\text{MPE}}$, both in expectation and in probability, and we’ve already established the same for the ADE above. Thus, by (47), $\tau_{\text{AIE}}$ must also converge to $\tau^{*}_{\text{AIE}}=\tau^{*}_{\text{MPE}}-\tau^{*}_{\text{ADE}}$, both in expectation (Theorem 2) and in probability (Theorem 4).
Lemma 15.
Under the setting of Theorem 2, suppose furthermore that the market may have random price perturbations $U_{i}$ as in Design 2 with $\mathbb{E}\left[U_{i}\right]=0$ and $\left\lVert U_{i}\right\rVert\leq h_{n}$ almost surely with $h_{n}=o(n^{-1/4})$. Then,
$\tau_{\text{ADE}}=\frac{1}{n}\sum\limits_{i=1}^{n}\left(Y_{i}(1,p^{*}_{\pi})-Y_{i}(0,p^{*}_{\pi})-\pi_{i}\Delta_{i}(1,\,p^{*}_{\pi})-(1-\pi_{i})\Delta_{i}(1,\,p^{*}_{\pi})\right)+o_{p}(1/\sqrt{n}),$ (48)
with $\Delta_{i}(w,\,p)$ as defined in (12). Furthermore, writing $\varepsilon_{i}(w)=Y_{i}(w,p^{*}_{\pi})-y(w,p^{*}_{\pi})$, we have
$\sqrt{n}\left(\tau_{\text{ADE}}-\tau^{*}_{\text{ADE}}\right)\Rightarrow\mathcal{N}\left(0,\,\operatorname{Var}\left[\varepsilon_{i}(1)-\varepsilon_{i}(0)-\pi_{i}\Delta_{i}(1,\,p^{*}_{\pi})-(1-\pi_{i})\Delta_{i}(1,\,p^{*}_{\pi})\right]\right).$ (49)
Lemma 16.
Under the setting of Lemma 15,
$\begin{split}&\mathbb{E}\left[\left\lvert\tau_{\text{MPE}}-\tau^{*}_{\text{MPE}}\right\rvert\right]=o(1),\\ &\tau^{*}_{\text{MPE}}=y(1,\,p^{*}_{\pi})-y(0,\,p^{*}_{\pi})+\xi_{y}^{\top}\xi_{z}^{-1}\left(z(1,\,p^{*}_{\pi})-z(0,\,p^{*}_{\pi})\right).\end{split}$ (50)
Proof of Lemma 15.
By definition of the ADE, we have
$\tau_{\text{ADE}}$ $=\mathbb{E}_{\bm{\pi}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,P_{n}(w_{i}=1;\,\bm{W}_{-i})+U_{i})-Y_{i}(0,P_{n}(w_{i}=0;\,\bm{W}_{-i})+U_{i})\right)\right]$
$=\mathbb{E}_{\bm{\pi}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,P_{n}(\bm{W})+U_{i})-Y_{i}(0,\,P_{n}(\bm{W})+U_{i})\right)+R_{n}(1)-R_{n}(0)\right],$
$R_{n}(w)$ $=\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(w,\,P_{n}(w_{i}=w;\,\bm{W}_{-i})+U_{i})-Y_{i}(w,\,P_{n}(\bm{W})+U_{i})\right).$
[…]
$\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(w,\,P_{n}(\bm{W})+U_{i})-y_{n}(w,\,P_{n}(\bm{W}))\right)$
$\quad\quad\quad\quad=\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(w,\,p^{*}_{\pi}+U_{i})-y_{n}(w,\,p^{*}_{\pi})\right)+o_{p}(1/\sqrt{n}),$
with $y_{n}(w,\,p)=\mathbb{E}\left[Y_{i}(w,\,p+U_{i})\right]$. Furthermore, arguing via twice differentiability of $y(w,\ p)$ like in the proof of Lemma 14, and noting the established rate of convergence $\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert_{2}=\mathcal{O}_{p}(1/\sqrt{n})$ in that result, we get
$\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(w,\,P_{n}(\bm{W})+U_{i})-Y_{i}(w,\,p^{*}_{\pi}+U_{i})\right)=\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\cdot\nabla_{p}y(w,\,p^{*}_{\pi})+o_{p}(1/\sqrt{n}).$
Combining our results so far, we get
$\tau_{\text{ADE}}=\mathbb{E}_{\bm{\pi}}\Bigg{[}\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,p^{*}_{\pi}+U_{i})-Y_{i}(0,\,p^{*}_{\pi}+U_{i})\right)$
$\quad\quad\quad\quad\quad\quad+\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\cdot\nabla_{p}\left(y(1,\,p^{*}_{\pi})-y(0,\,p^{*}_{\pi})\right)\Bigg{]}+o_{p}(1/\sqrt{n}),$
where we note that the $o_{p}(1/\sqrt{n})$ residual term can be moved outside the $\mathbb{E}_{\pi}\left[.\right]$ term because it is almost surely bounded. Finally, plugging in the asymptotic expansion (40) for $P_{n}(\bm{W})-p^{*}_{\pi}$ into this expression gives
$\tau_{\text{ADE}}=\mathbb{E}_{\bm{\pi}}\Bigg{[}\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,p^{*}_{\pi}+U_{i})-Y_{i}(0,\,p^{*}_{\pi}+U_{i})\right)$
$\quad\quad\quad\quad\quad\quad+\nabla_{p}^{\top}\left(y(1,\,p^{*}_{\pi})-y(0,\,p^{*}_{\pi})\right)\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\Bigg{]}+o_{p}(1/\sqrt{n}),$
which in turn yields (48) by integrating over $W_{i}$. Finally, the central limit theorem follows by an analogous argument to the one made in Lemma 14.
We now return to arguing that the residual terms $R_{n}(w)$ are small. Thanks to Assumption 4, there exists a constant $C$ such that
$\left\lvert R_{n}(w)\right\rvert$ $\leq\frac{1}{n}\sum_{i=1}^{n}\left\lvert Y_{i}(w,\,P_{n}(w_{i}=w;\,\bm{W}_{-i})+U_{i})-Y_{i}(w,\,P_{n}(\bm{W}_{-i})+U_{i})\right\rvert$
$\leq\frac{1}{n}\sum_{i=1}^{n}\left\lvert Y_{i}(w,\,P_{n}(w_{i}=1;\,\bm{W}_{-i})+U_{i})-Y_{i}(w,\,P_{n}(w_{i}=0;\,\bm{W}_{-i})+U_{i})\right\rvert$
$\leq C\,\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{J}\left\lvert Z_{ij}(w,\,P_{n}(w_{i}=1;\,\bm{W}_{-i})+U_{i})-Z_{ij}(w,\,P_{n}(w_{i}=0;\,\bm{W}_{-i})+U_{i})\right\rvert$
$\quad\quad\quad\quad+C\,\frac{1}{n}\sum_{i=1}^{n}\left\lVert P_{n}(w_{i}=1;\,\bm{W}_{-i})-P_{n}(w_{i}=0;\,\bm{W}_{-i})\right\rVert_{2}.$
We can now invoke Lemma 17 as well as approximate monotonicity as in Assumption 3 to check that, for any $\varepsilon>0$,
$\left\lvert R_{n}(w)\right\rvert$ $\leq C\,\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{J}\Big{(}Z_{ij}(w,\,P_{n}(\bm{W})-\varepsilon/\sqrt{n}e_{j}+U_{i})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-Z_{ij}(w,\,P_{n}(\bm{W})+\varepsilon/\sqrt{n}e_{j}+U_{i})\Big{)}+o_{p}(1/\sqrt{n})$
$\leq-\frac{2C\varepsilon}{\sqrt{n}}\sum_{j=1}^{J}\frac{\partial}{\partial p_{j}}z_{j}(w,\ p^{*}_{\pi})+o_{p}(1/\sqrt{n}),$
where the second line followed from asymptotic equicontinuity as used in the proof of Lemma 15. Because this bound holds for any $\varepsilon>0$, we in fact have $R_{n}(w)=o_{p}(1/\sqrt{n})$. ∎
Lemma 17.
Under the conditions of Lemma 14, changing the treatment given to any one unit cannot change the equilibrium price much:
$\sup_{i=1,\ldots,\ n}\left\lVert P_{n}(w_{i}=1;\bm{W}_{-i})-P_{n}(w_{i}=0;\bm{W}_{-i})\right\rVert_{2}=o_{p}(1/\sqrt{n}).$ (51)
Proof.
By Assumption 1 and a union bound, we know that with probability at least $1-2ne^{-cn}$,
$\left\lVert\frac{1}{n}\left(Z_{i}\left(w,\,P_{n}(w_{i}=w,\,\bm{W}_{-i})+U_{i}\right)+\sum_{j\neq i}Z_{j}\left(W_{j},\,P_{n}(w_{i}=w,\,\bm{W}_{-i})+U_{j}\right)\right)\right\rVert_{2}\leq a_{n}$ (52)
[…]
$\left\lVert\frac{1}{n}\sum_{j=1}^{n}Z_{j}\left(W_{j},\,P_{n}(w_{i}=w,\,\bm{W}_{-i})+U_{j}\right)\right\rVert_{2}\leq a_{n}+\frac{2M}{n}=o(1/\sqrt{n}).$ (53)
We can thus use asymptotic equicontinuity as in the proof of Lemma 15 to argue that
$\sup_{i,w}\left\lvert P_{n}(w_{i}=w,\,\bm{W}_{-i})-p^{*}_{\pi}+\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right\rvert=o_{p}(1/\sqrt{n}),$
which implies our desired result as the asymptotic linear representation is the same regardless of $i$ and $w$. ∎
Proof of Lemma 16.
Recall that $\tau_{\text{MPE}}$ is defined as
$\tau_{\text{MPE}}=\sum_{k=1}^{n}\frac{\partial}{\partial\pi_{k}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\pi}\left[Y_{i}(W_{i},P_{n}(\bm{W})+U_{i}\right].$
Writing $f_{\pi}(\bm{w})$ for the distribution function of $\bm{w}$, we can write this more explicitly as
$\tau_{\text{MPE}}=\sum\limits_{k=1}^{n}\frac{\partial}{\partial\pi_{k}}\sum\limits_{\bm{w}}f(\bm{w})\,\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(w_{i},P_{n}(\bm{W})+U_{i}),$
Now, under our design, $W_{i}\sim\mbox{Bernoulli}(\pi_{i})$ independently for each sample, and so $f_{\pi}(\bm{w})=\prod_{k=1}^{n}\pi_{k}^{w_{k}}(1-\pi_{k})^{1-w_{k}}$. Taking the derivative with respect to $\pi_{k}$ yields
$\frac{\partial}{\partial\pi_{k}}f_{\pi}(\bm{w})=\frac{w_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}f(\bm{w}),$
and so
$\begin{split}\tau_{\text{MPE}}&=\sum\limits_{k=1}^{n}\sum\limits_{\bm{w}}\frac{\partial}{\partial\pi_{k}}f(\bm{w})\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(w_{i},P_{n}(\bm{W})+U_{i})\\ &=\mathbb{E}_{\pi}\left[\sum\limits_{k=1}^{n}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},P_{n}(\bm{W})+U_{i})\right].\end{split}$ (54)
At the end of this proof, we will show that
$\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},P_{n}(\bm{W})+U_{i})=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+U_{i})-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+B_{n},$ (55)
where the remainder term $B_{n}$ satisfies $\lim_{n\rightarrow\infty}n\mathbb{E}\left[B_{n}^{2}\right]$ = 0. By Cauchy-Schwarz
$\mathbb{E}\left[\left\lvert\mathbb{E}_{\pi}\left[\sum\limits_{k=1}^{n}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right]\right\rvert\right]$ $\leq\mathbb{E}\left[\mathbb{E}_{\pi}\left[\sum\limits_{k=1}^{n}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right]^{2}\right]$
$\leq\mathbb{E}\left[n\mathbb{E}_{\pi}\left[B_{n}^{2}\right]\right]=n\mathbb{E}\left[B_{n}^{2}\right]=o(1),$
and so this remainder term in fact has a lower-order effect on $\tau_{\text{MPE}}$.
[…]
$\mathbb{E}_{\pi}\left[\sum\limits_{k=1}^{n}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+U_{i})\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\left(Y_{i}(1,p^{*}_{\pi}+U_{i})-Y_{i}(0,p^{*}_{\pi}+U_{i})\right)$
$\mathbb{E}_{\pi}\left[\sum\limits_{k=1}^{n}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right]=\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}\left(Z_{i}(1,p^{*}_{\pi}+U_{i})-Z_{i}(0,p^{*}_{\pi}+U_{i})\right),$
thus immediately implying the desired result (as these quantities are IID averages that concentrate around the claimed terms in $\tau^{*}_{\text{MPE}}$).
We now return to verifying that $\mathbb{E}\left[B_{n}^{2}\right]=o(1/n)$. We can expand out this residual as
$B_{n}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},P_{n}(\bm{W})+U_{i})-\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+U_{i})-\left(y_{n}\left(P_{n}(\bm{W})+U_{i})\right)-y_{n}\left(p^{*}_{\pi}\right)\right)$
$\quad\quad+y_{n}\left(P_{n}(\bm{W})+U_{i})\right)-y_{n}\left(p^{*}_{\pi}\right)-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i}).$
Furthermore, (43) in the proof of Lemma 14 gives us a tail bound for $\sqrt{n}\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert_{2}$ with an exponential tail decay. In particular, $\mathbb{P}\left[\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert_{2}>n^{-1/3}\right]$ decays exponentially fast in $n$ (because $n^{-1/3}=t_{n}/\sqrt{n}$ for $t_{n}=n^{1/6}$) so we can use the $L_{2}$-bound on the residual in Lemma 20 to verify that the second moment of the asymptotic equicontinuity residual on the first line of our decomposition decays as $o(1/n)$. Meanwhile,
$y_{n}\left(P_{n}(\bm{W})+U_{i})\right)-y_{n}\left(p^{*}_{\pi}\right)-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})$
$\quad\quad\quad\quad=y\left(P_{n}(\bm{W})+U_{i})\right)-y\left(p^{*}_{\pi}\right)-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+o(1/\sqrt{n}),$
and
$\left(P_{n}(\bm{W})+U_{i})-p^{*}_{\pi}\right)y^{\prime}\left(p^{*}_{\pi}\right)=\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+o(1/\sqrt{n}).$
Thus
$\mathbb{E}\left[\left(y_{n}\left(P_{n}(\bm{W})+U_{i})\right)-y_{n}\left(p^{*}_{\pi}\right)-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right)^{2}\right]$
$\quad\quad\quad\quad=\mathbb{E}\left[\left(y\left(P_{n}(\bm{W})+U_{i})\right)-y\left(p^{*}_{\pi}\right)-\left(P_{n}(\bm{W})+U_{i})-p^{*}_{\pi}\right)y^{\prime}\left(p^{*}_{\pi}\right)\right)^{2}\right]+o(1/n)$
$\quad\quad\quad\quad=\mathcal{O}\left(\mathbb{E}\left[\left\lVert P_{n}(\bm{W})+U_{i})-p^{*}_{\pi}\right\rVert_{2}^{2}\right]\right)+o(1/n)=o(1/n),$
where on the last line we used continuous twice differentiability of $y(p)$ and the bound (43) to control the second moment of market price fluctuations. ∎
B.3 Proof of Theorem 5
Thanks to the self-normalized form of the difference-in-means estimator, we have
$\hat{\tau}_{\text{ADE}}-\tau^{*}_{\text{ADE}}=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\left(Y_{i}-y(1,\,p^{*}_{\pi})\right)}{\hat{\pi}}-\frac{(1-W_{i})\left(Y_{i}-y(1,\,p^{*}_{\pi})\right)}{1-\hat{\pi}}.$ (56)
By asymptotic equicontinuity as shown in Lemma 20 followed by an application of Theorem 1,
$\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\left(Y_{i}-y(1,p^{*}_{\pi})\right)}{\pi}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\,\varepsilon_{i}(1)}{\pi}+y(1,\,P_{n})-y(1,\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\,\varepsilon_{i}(1)}{\pi}-\nabla_{p}^{\top}y(1,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right),$
$\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\left(Y_{i}-y(0,p^{*}_{\pi})\right)}{1-\pi}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\,\varepsilon_{i}(1)}{1-\pi}+y(0,\,P_{n})-y(0,\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\,\varepsilon_{i}(1)}{1-\pi}-\nabla_{p}^{\top}y(0,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right),$
and furthermore a central limit theorem follows by standard Lindeberg arguments,
$\frac{1}{\sqrt{n}}\begin{pmatrix}\sum_{i=1}^{n}\frac{W_{i}\left(Y_{i}-y(1,p^{*}_{\pi})\right)}{\pi}\\ \sum_{i=1}^{n}\frac{(1-W_{i})\left(Y_{i}-y(0,p^{*}_{\pi})\right)}{1-\pi}\end{pmatrix}\Rightarrow\left(\begin{pmatrix}0\\ 0\end{pmatrix},\ \operatorname{Cov}\left[\begin{pmatrix}\frac{W_{i}\,\varepsilon_{i}(1)}{\pi}-\nabla_{p}^{\top}y(1,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\\ \frac{(1-W_{i})\,\varepsilon_{i}(0)}{1-\pi}-\nabla_{p}^{\top}y(0,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\end{pmatrix}\right]\right).$
The result then follows from Slutsky’s lemma because $\hat{\pi}\rightarrow_{p}\pi$.
B.4 Proof of Proposition 6
Throughout this proof, all quantities are computed in population equilibrium and so we use the short-hand $\Delta_{i}(w)=\Delta_{i}(w,\,p^{*}_{\pi})$. Expanding the square, we can verify that
$\sigma^{2}_{D}=\mathbb{E}\left[\frac{\varepsilon_{i}^{2}(1)}{\pi}+\frac{\varepsilon_{i}^{2}(0)}{1-\pi}\right]-2\mathbb{E}\left[\varepsilon_{i}(1)\Delta_{i}(1)-\varepsilon_{i}(0)\Delta_{i}(0)\right]+\mathbb{E}\left[\pi\Delta_{i}^{2}(1)+(1-\pi)\Delta_{i}^{2}(0)\right].$
Similarly, we have
$\bar{\sigma}^{2}_{D}$ $=\mathbb{E}\left[\left(\frac{1}{\pi}-1\right)\varepsilon_{i}^{2}(1)+\left(\frac{1}{1-\pi}-1\right)\varepsilon_{i}^{2}(0)+2\varepsilon_{i}(1)\varepsilon_{i}(0)\right]$
$\quad\quad-2\mathbb{E}\left[\left((1-\pi)\varepsilon_{i}(1)+\pi\varepsilon_{i}(0)\right)\left(\Delta_{i}(1)-\Delta_{i}(0)\right)\right]$
$\quad\quad+\mathbb{E}\left[\pi(1-\pi)\left(\Delta_{i}^{2}(1)-2\Delta_{i}(1)\Delta_{i}(0)+\Delta_{i}^{2}(0)\right)\right].$
Taking the difference and unexpanding squares where possible, we get
$\sigma^{2}_{D}-\bar{\sigma}^{2}_{D}$ $=\mathbb{E}\left[\left(\varepsilon_{i}(1)-\varepsilon_{i}(0)\right)^{2}\right]-2\mathbb{E}\left[\left(\varepsilon_{i}(1)-\varepsilon_{i}(0)\right)\left(\pi\Delta_{i}(1)+(1-\pi)\Delta_{i}(0)\right)\right]$
$\quad\quad\quad\quad\quad+\mathbb{E}\left[\left(\pi\Delta_{i}(1)+(1-\pi)\Delta_{i}(0)\right)^{2}\right]$
$=\mathbb{E}\left[\left(\varepsilon_{i}(1)-\varepsilon_{i}(0)-\left(\pi\Delta_{i}(1)+(1-\pi)\Delta_{i}(0)\right)\right)^{2}\right]$
$\geq 0,$
as claimed.
B.5 Proof of Theorem 7
For simplicity, defined the unscaled perturbation $S_{i}=U_{i}/h_{n}$, and recall that $S_{i}\in\left\{\pm 1\right\}^{J}$ uniformly at random. Our estimator of the average indirect effect can equivalently be written as
$\hat{\tau}_{\text{AIE}}=-(\mathbf{Y}^{\top}\mathbf{S})\left(\mathbf{Z}^{\top}\mathbf{S}\right)^{-1}\hat{\tau}_{\text{ADE}}^{z}.$
One can replicate the proof of Theorem 5 to verify that $\hat{\tau}_{\text{ADE}}^{z}=z(1,\,p^{*}_{\pi})-z(0,\,p^{*}_{\pi})+\mathcal{O}_{P}(1/\sqrt{n})$, and thus makes an asymptotically negligible contribution to the error.
Next, by asymptotic equicontinuity as in Lemma 20,
$\frac{1}{n}\mathbf{Y}^{\top}\mathbf{S}$ $=\frac{1}{n}\sum_{i=1}^{n}Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})S_{i}+\frac{h_{n}}{2^{J}}\sum_{v\in\left\{\pm 1\right\}^{J}}\left(y(P_{n}+h_{n}v)-y(p^{*}_{\pi}+h_{n}v)\right)v+o_{p}(1/\sqrt{n})$
$=\frac{1}{n}\sum_{i=1}^{n}Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})S_{i}+\mathcal{O}(h_{n}^{2})+o_{p}(1/\sqrt{n})$
by twice continuous differentiability of $y$; a similar expansion holds for $\mathbf{Z}^{\top}\mathbf{S}$. Because $h_{n}=o(1/\sqrt[4]{n})$, we can then fold the $\mathcal{O}(h_{n}^{2})$ term into the $o_{p}(1/\sqrt{n})$ error. Again by twice differentiability of $y(\cdot)$, we can invoke Taylor expansion to verify that
$\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}$
[…]
$\quad\quad\quad\quad\quad\quad+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n}),$
$\quad\quad=h_{n}\xi_{y}+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})+y(p^{*}_{\pi})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n}),$
and a similar decomposition holds for $\text{vec}\left(Z_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}$. Furthermore
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})-\left(Y_{i}(W_{i},\,p^{*}_{\pi})-y(p^{*}_{\pi})\right)\right)^{2}=0$
by weak continuity as in Assumptions 5 and 4, and these error terms are all mean-zero conditionally on $S_{i}$ and uncorrelated. Thus,
$\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}=h_{n}\xi^{\top}_{y}+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n});$
a similar expansion again holds for the $Z_{i}$ term.
Putting everything together, and writing $\mathbf{Y}^{*}$ for the vector with $i$-th entry $Y_{i}(W_{i},\,p^{*}_{\pi})$ and $\mathbf{Z}^{*}$ for the matrix with $i$-th row $Z_{i}(W_{i},\,p^{*}_{\pi})$, we then get
$(\mathbf{Y}^{\top}\mathbf{S})\left(\mathbf{Z}^{\top}\mathbf{S}\right)^{-1}$
$\quad=(\mathbf{Y}^{\top}\mathbf{S})\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\left(\mathbf{Z}^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\right)^{-1}$
$\quad=\left(h_{n}\xi^{\top}_{y}+\mathbf{Y}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\right)\left(h_{n}\xi_{z}+\mathbf{Z}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\right)^{-1}+o_{p}\left(\frac{h_{n}}{\sqrt{n}}\right)$
$\quad=\xi^{\top}_{y}\xi_{z}^{-1}+h_{n}^{-1}\mathbf{Y}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}-h_{n}^{-1}\xi^{\top}_{y}\xi_{z}^{-1}\mathbf{Z}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}+o_{p}\left(\frac{h_{n}}{\sqrt{n}}\right),$
where the last step is by Taylor expansion. Finally, $n\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}\rightarrow_{p}\xi_{z}^{-1}$, and so by Slutsky’s lemma
$\hat{\tau}_{\text{AIE}}-\tau^{*}_{\text{AIE}}=-\frac{1}{nh_{n}}\sum_{i=1}^{n}\left(Y_{i}(W_{i},\,p^{*}_{\pi})-\xi_{y}^{\top}\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\right)S_{i}^{\top}\xi_{z}^{-1}\tau^{*}_{ADE,Z}+o_{p}\left(\frac{h_{n}}{\sqrt{n}}\right).$
Plugging in $U_{i}=h_{n}S_{i}$ then yields the desired asymptotic linear expansion. The central limit theorem again follows by a standard Lindeberg argument.
B.6 Proof of Theorem 8
All parameter estimates plugged into (23) are consistent, and so $\hat{\sigma}^{2}_{D}=\hat{\sigma}^{*2}_{D}+o_{p}(1)$ with
$\hat{\sigma}^{*2}_{D}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}\left(Y_{i}-y(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n})\right)^{2}.$ (57)
If $\hat{\sigma}^{*2}_{D}$ were an IID average, the claim from the theorem statement would be immediate; however, we still need to address dependence due to the shared price shocks from $P_{n}$.
To this end, we introduce an auxiliary random variable $T_{i}\in\left\{0,2\right\}$ uniformly at random. It is well known from the literature on resampling-based variance estimation (e.g., Præstgaard & Wellner 1993) that
$\begin{split}&\hat{\sigma}^{*2}_{D}=n\operatorname{Var}_{T}\left[\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}-y(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n})\right)\right],\end{split}$ (58)
where $\operatorname{Var}_{T}\left[.\right]$ denotes variance induced by only the random weights $T_{i}$. Now, by asymptotic equicontinuity as shown in Lemma 20,
$\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}-y(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n})\right)$
$\quad\quad=\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}(1,\,p^{*}_{\pi})-y(1,\,p^{*}_{\pi})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}(0,\,p^{*}_{\pi})-y(0,\,p^{*}_{\pi})\right)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right)$
$\quad\quad\quad\quad\quad\quad+\nabla_{p}^{\top}[y(1,\,p^{*}_{\pi})-y(0,\,p^{*}_{\pi})]\xi_{z}^{-1}\left(z\left(P_{n}\right)-z\left(p^{*}_{\pi}\right)\right)+B_{n}(\mathbf{T}),$
with $\lim_{n\rightarrow\infty}n\mathbb{E}\left[B_{n}(\mathbf{T})^{2}\right]=0$. Thus, from (58), we immediately see that
$\begin{split}\hat{\sigma}^{*2}_{D}&=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right)^{2}+n\operatorname{Var}_{T}\left[B_{n}(\mathbf{T})\right]\\ &\quad\quad\quad\quad+2\operatorname{Cov}_{T}\left[\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right),\,B_{n}(\mathbf{T})\right],\end{split}$ (59)
with notation $\varepsilon_{i}(w)$ as defined in the statement of Theorem 5. Now,
$n\mathbb{E}\left[\operatorname{Var}_{T}\left[B_{n}(\mathbf{T})\right]\right]\leq n\operatorname{Var}\left[B_{n}(\mathbf{T})\right]\leq n\mathbb{E}\left[B_{n}(\mathbf{T})^{2}\right]=o(1),$ (60)
and so the terms depending on $B_{n}(\mathbf{T})$ are asymptotically negligible; the covariance term needs to be bounded by Cauchy-Schwarz before applying (60). Finally, the dominant term in (59) is an IID average whose mean is $\sigma^{2}_{D}$, and thus our result follows by standard concentration arguments.
The consistency of $\hat{\sigma}_{I}^{2}$ can be established via an analogous argument, which is omitted here for conciseness.
B.7 Proof of Proposition 9
The proof of Theorem 1 in Appendix B.1 of Hu et al. (2022) shows that under a general interference pattern that $\frac{\partial}{\partial\pi^{\prime}_{k}}\mathbb{E}_{\pi^{\prime}}[Y_{i}(\bm{W})]_{\pi^{\prime}=\pi}=\mathbb{E}_{\pi}\left[Y_{i}(w_{k}=1;W_{-k})-Y_{i}(w_{k}=0;W_{-k})\right].$ Plugging this into our conditional expectation of interest,
$\mathbb{E}\left[\frac{\partial}{\partial\pi_{k}}\sum\limits_{i=1}^{n}\mathbb{E}_{\pi}[Y_{i}(\bm{W})]\Big{|}X_{k}=x\right]$ $=\mathbb{E}\left[\sum\limits_{i=1}^{n}\mathbb{E}_{\pi}\left[Y_{i}(w_{k}=1;W_{-k})-Y_{i}(w_{k}=0;W_{-k})\right]\Big{|}X_{k}=x\right]$
$=\sum\limits_{i=1}^{n}\mathbb{E}\left[Y_{i}(w_{k}=1;W_{-k})-Y_{i}(w_{k}=0;W_{-k})|X_{k}=x\right]$
$=\mathbb{E}\left[Y_{i}(w_{i}=1;W_{-k})-Y_{i}(w_{i}=0;W_{-i})|X_{i}=x\right]$
$\qquad+\sum\limits_{j\neq i}\mathbb{E}\left[Y_{j}(w_{i}=1;W_{-i})-Y_{j}(w_{i}=0;W_{-i})|X_{i}=x\right]$
$=\bar{\tau}_{\text{CADE}}(x)+\bar{\tau}_{\text{CAIE}}(x)$
B.8 Proof of Theorem 10
[…]
$\left\lvert Y_{i}\left(w,\,P(w_{i}=w,\,W_{-i})\right)-Y_{i}\left(w,\,p^{*}_{\pi}\right)\right\rvert$
$\quad\quad\quad\quad\leq M\left(\delta_{P}(w)+\sum_{j=1}^{J}\left(Z_{ij}\left(w,\,p^{*}_{\pi}-C\delta_{P}(w)e_{j}\right)-Z_{ij}\left(w,\,p^{*}_{\pi}+C\delta_{P}(w)e_{j}\right)\right)\right),$
[…]
$\mathbb{P}\left[\left\lVert P(w_{i}=w,\,W_{-i})-p^{*}_{\pi}\right\rVert_{2}>n^{-1/3}\,\big{|}\,X_{i}=x\right]=\mathcal{O}\left(e^{-n^{1/3}}\right).$
Thus, thanks to weak continuity of net demand as in Assumption 5,
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\left\lvert\mathbb{E}_{\pi}\left[Y_{i}\left(w,\,P(w_{i}=w,\,W_{-i})\right)\right]-Y_{i}\left(w,\,p^{*}_{\pi}\right)\right\rvert\,\big{|}\,X_{i}=x\right]=0.$
We can then conclude that
$\lim_{n\rightarrow\infty}\tau_{\text{CADE}}(x)-\mathbb{E}\left[Y_{i}\left(1,\,p^{*}_{\pi}\right)-Y_{i}\left(0,\,p^{*}_{\pi}\right)\,\big{|}\,X_{i}=x\right]=0,$
which implies the first part of the claimed result.
For the second claim, we follow the structure from the proof of Theorem 2 and note that, by Proposition 9, it suffices to show that
$\mathbb{E}\left[\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\,\big{|}\,X_{k}\in S\right]=\mathbb{E}\left[\tau^{*}_{\text{CADE}}(x)+\tau^{*}_{\text{CAIE}}(x)\,\big{|}\,X_{k}\in S\right].$ (61)
We have chosen the set $S$ such that $\mathbb{P}\left[X_{k}\in S\right]>0$. Thus, there is a $c>0$ such that with probability $1-e^{cn}$ there is at least one observation $X_{k}\in S$, and so, using a convention that $/0=0$ 0,
$\mathbb{E}\left[\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\,\big{|}\,X_{k}\in S\right]=\mathbb{E}\left[\frac{1}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\right]+\mathcal{O}\left(e^{-cn}\right).$
We can then follow the proof of Lemma 15, including the transformation of a derivative into a covariance as in (54) and the expansion (55), to verify that
$\begin{split}&\mathbb{E}\left[\frac{1}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\right]\\ &\quad\quad\quad=\mathbb{E}\left[\tau^{*}_{\text{CADE}}(x)+\tau^{*}_{\text{CAIE}}(x)\,\big{|}\,X_{k}\in S\right]\mathbb{P}\left[\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert>0\right]\\ &\quad\quad\quad\quad\quad\quad+\mathbb{E}\left[\frac{n}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right],\end{split}$ (62)
with $\lim_{n\rightarrow\infty}n\mathbb{E}\left[B_{n}^{2}\right]=0$. We have that
$\mathbb{E}\left[\left\lvert\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right\rvert\right]=0$
by Cauchy-Schwarz as in the proof of Lemma 15, and that the term inside the expectation admits an almost sure bound that scales polynomially in $n$. Meanwhile, $n\,/\,\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert\leq 2/\mathbb{P}\left[X_{k}\in S\right]$ with probability that decays exponentially in $n$. Thus, the whole expectation in (62) goes to 0 with $n$, as desired.
B.9 Proof of Theorem 11
By treatment randomization |$\left\lvert\left\{i\in N_{k}(x):W_{i}=1\right\}\right\rvert/k\rightarrow_{p}\pi$, and so
$\hat{\tau}_{\text{CADE}}(x)=\frac{1}{k}\sum_{\left\{i\in N_{k}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}-\frac{1-W_{i}}{1-\pi}\,Y_{i}\right).$ (63)
Now, let $B_{\varepsilon}(x)=\left\{x^{\prime}:\left\lVert x-x^{\prime}\right\rVert_{2}\leq\varepsilon\right\}$. For any $\varepsilon>0$, we have $\mathbb{P}\left[X\in B_{\varepsilon}(x)\right]>0$ and so, because $k/n\rightarrow 0$, the $k$-nearest neighbor set $N_{k}(x)$ is asymptotically concentrated near $x$, i.e., $\lim_{n\rightarrow\infty}\mathbb{P}\left[\left\{X_{i}:i\in N_{k}(x)\right\}\subset B_{\varepsilon}(x)\right]=1$ for any $\varepsilon>0$. Thus, because the conditional distribution of $Y_{i}(w,\,p)$ given $X_{i}=x$ varies continuously in $x$ and $Y_{i}(w,\,p)$ is bounded, conditional moments of $Y_{i}(w,\,p)$ also vary continuously in $x$ and so we can follow the proof of Lemma 20 to verify that, for any sequence $p_{n}\rightarrow_{p}p$,
$\begin{split}\lim_{n\rightarrow\infty}\left\lvert\frac{1}{k}\sum_{N_{k}(x)}\frac{W_{i}}{\pi}\,Y_{i}(1,p_{n})-\frac{1}{k}\sum_{N_{k}(x)}\left(\frac{W_{i}}{\pi}\,Y_{i}(1,p)+y(1,p_{n};\,X_{i})-y(1,p;\,X_{i})\right)\right\rvert=o_{p}\left(\frac{1}{\sqrt{k}}\right),\end{split}$ (64)
and a similar expansion holds for $(1-W_{i})Y_{i}(0,\,p_{n})/(1-\pi)$. Because $P_{n}\rightarrow p^{*}_{\pi}$ by Theorem 1 and $y(w,\,p;\,x)$ is continuous in $p$ and $x$ (and thus uniformly continuous in $p$ in a neighborhood of $x$), we have that
$\frac{1}{k}\sum_{\left\{i\in N_{k}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}-\frac{1-W_{i}}{1-\pi}\,Y_{i}\right)=\frac{1}{k}\sum_{\left\{i\in N_{k}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}(1,\,p^{*}_{\pi})-\frac{1-W_{i}}{1-\pi}\,Y_{i}(0,\,p^{*}_{\pi})\right)+o_{p}(1).$ (65)
The result then follows by concentration because $k\rightarrow\infty$. We note that this argument can be extended to obtain rates of convergence under smoothness assumptions following Stone (1980).
B.10 Proof of Proposition 12
By uniqueness of equilibrium prices, $p^{*}_{\nu}=p^{*}_{\pi}$ if and only if
$\mathbb{E}[\nu(X_{i})Z_{i}(1,p^{*}_{\pi})+(1-\nu(X_{i}))Z_{i}(0,p^{*}_{\pi})]=\mathbb{E}[\pi(X_{i})Z_{i}(1,p^{*}_{\pi})+(1-\pi(X_{i}))Z_{i}(0,p^{*}_{\pi})]=0.$
This means the constraint $p^{*}_{\nu}=p^{*}_{\pi}$ in optimization problem in Equation (35) can be replaced with the above moment constraint. We can then transform the constraint by subtracting $\mathbb{E}[Z_{i}(0,p^{*}_{\pi})]$ from both sides above and applying the chain rule to get the constraint claimed in Proposition 12. Meanwhile, for the objective, we can use the equilibrium stability condition to verify
$\mathbb{E}\left[Y_{i}(W_{i},p^{*}_{\nu})\right]$ $=\mathbb{E}\left[\nu(X_{i})Y_{i}(1,\,p^{*}_{\nu})+(1-\nu(X_{i}))Y_{i}(0,\,p^{*}_{\nu})\right]$
$=\mathbb{E}\left[\nu(X_{i})Y_{i}(1,\,p^{*}_{\pi})+(1-\nu(X_{i}))Y_{i}(0,\,p^{*}_{\pi})\right]$
$=\mathbb{E}\left[\nu(X_{i})\left(Y_{i}(1,\,p^{*}_{\pi})-Y_{i}(0,\,p^{*}_{\pi})\right)\right]+\mathbb{E}\left[Y_{i}(0,\,p^{*}_{\pi})\right]$
$=\mathbb{E}\left[\nu(X_{i})\mathbb{E}\left[Y_{i}(1,\,p^{*}_{\pi})-Y_{i}(0,\,p^{*}_{\pi})\,\big{|}\,X_{i}\right]\right]+\mathbb{E}\left[Y_{i}(0,\,p^{*}_{\pi})\right],$
which yields the desired expression after dropping the last term that does not depend on $\nu(\cdot)$.
B.11 Proof of Theorem 13
For convenience, we use the following short-hand for the policy value and the asymptotic budget imbalance:
$V(\nu)=\mathbb{E}\left[\nu(X_{i})\tau_{\text{CADE}}^{*}(X_{i})\right],\ \ \ \ G(\nu)=\mathbb{E}\left[\left(\nu(X_{i})-\pi(X_{i})\right)\tau_{\text{CADE}}^{*,z}(X_{i})\right].$ (66)
Then, following the setup of Luenberger (1969, Chapter 8.6), we define primal and dual objectives over $\theta,\,\lambda\in\mathbb{R}_{+}^{J}$ as
$\omega(\theta)=\inf_{\nu}\left\{-V(\nu):\left\lvert G_{j}(\nu)\right\rvert\leq\theta_{j},\,j=1,\,\ldots,\,J\right\},\ \ \ \ \phi(\lambda)=\inf_{\nu}\left\{-V(\nu)+\lambda\cdot\left\lvert G(\nu)\right\rvert\right\},$ (67)
[…]
$\omega(\theta)=\sup\left\{\phi(\lambda):\lambda\geq\theta\right\}$ (68)
for all $\theta$ with $\min\left\{\theta_{j}\right\}>0$. The dual function $\phi$ is clearly continuous at 0, and so we also see that for any sequence $\theta$ converging to 0,
$\lim_{\theta\rightarrow 0,\,\theta>0}\omega(\theta)=\sup\left\{\phi(\lambda):\lambda\geq 0\right\}.$ (69)
At the end of this proof, we will further show that $\omega(\theta)$ is continuous at 0, thus implying that in fact
$\omega(0)=\sup\left\{\phi(\lambda):\lambda\geq 0\right\},$ (70)
where $\omega(0)$ is the negative of the optimal value in our original optimization problem.
Now, given strong duality as in (70), we can express the policy value in our original problem as
$\begin{split}V(\nu^{*})&=\inf_{\lambda\in\mathbb{R}^{J}_{+}}\left\{\sup_{\nu}\left\{V(\nu)-\lambda\cdot\left\lvert G(\nu)\right\rvert\right\}\right\}.\end{split}$ (71)
Writing $\nu^{*}_{\lambda}$ for the solution of the inner optimization problem we see immediately from its first-order conditions that, for all $x$, one of the following 3 conditions must hold:
$\tau^{*}_{\text{CADE}}(x)-\sum_{j=1}^{J}\lambda_{j}\operatorname{sign}\left(G_{j}(\nu_{\lambda}^{*})\right)\tau^{*,z}_{CADE,j}(x)>0\ \text{ and}\ \nu_{\lambda}^{*}(x)=1,$
$\tau^{*}_{\text{CADE}}(x)-\sum_{j=1}^{J}\lambda_{j}\operatorname{sign}\left(G_{j}(\nu_{\lambda}^{*})\right)\tau^{*,z}_{CADE,j}(x)<0\ \text{ and}\ \nu_{\lambda}^{*}(x)=0,$
$\tau^{*}_{\text{CADE}}(x)-\sum_{j=1}^{J}\lambda_{j}\operatorname{sign}\left(G_{j}(\nu_{\lambda}^{*})\right)\tau^{*,z}_{CADE,j}(x)=0.$
Thus, equivalently, using notation from (37) we must have
[…]
Plugging back into (71), we get
[…]
Write $c^{*},\,a^{*}$ for the solution of the above optimization problem. We have
$V(\nu^{*})=V(\nu_{c^{*},a^{*}}(x))-\left\lvert c^{*}\right\rvert\cdot\left\lvert G(\nu_{c^{*},a^{*}}(x))\right\rvert\leq V(\nu_{c^{*},a^{*}}(x)),$
and so $V(\nu^{*})=V(\nu_{c^{*},a^{*}}(x))$ by optimality of the policy $\nu^{*}$.
We now return to verifying continuity of $\omega(\theta)$ at 0. Let
=$\mathcal{B}=\left\{b:\mathbb{E}\left[\left(\nu(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]=b\text{ for some }0\leq\nu(\cdot)\leq 1\right\}.$
[…]
$-\eta\mathcal{B}\subset\mathcal{B}.$ (72)
For simplicity, [^7] If the set doesn’t span $\mathbb{R}^{J}$ we can simply ignore the direction without any variation at the cost of extra notation. assume that $\mathcal{B}$ spans $\mathbb{R}^{J}$; thus, by (72) and convexity of $\mathcal{B}$, we know that there is a $\varepsilon>0$ such that $b\in\mathcal{B}$ for all $\left\lVert b\right\rVert_{2}\leq\varepsilon$.
Now, let $\omega(\theta)$ denote the primal objective (67) for some $\left\lVert\theta\right\rVert_{2}\leq\varepsilon$, and let $\nu^{*}_{\theta}$ be a maximizer of this objective. By the above argument, we know that there exists another policy $\nu^{\prime}_{\theta}$ with
$\mathbb{E}\left[\left(\nu^{\prime}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]=-\varepsilon\ \frac{\mathbb{E}\left[\left(\nu^{*}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]}{\left\lVert\mathbb{E}\left[\left(\nu^{*}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]\right\rVert_{2}}.$
We can construct a third policy by convex averaging of $\nu^{*}_{\theta}$ and $\nu^{\prime}_{\theta}$:
$\nu^{+}_{\theta}=\frac{\left\lVert\mathbb{E}\left[\left(\nu^{*}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]\right\rVert_{2}\nu^{\prime}_{\theta}+\varepsilon\nu^{*}_{\theta}}{\left\lVert\mathbb{E}\left[\left(\nu^{*}_{\theta}(X)-\pi(X)\right)\tau^{*,z}_{\text{CADE}}(X)\right]\right\rVert_{2}+\varepsilon}.$
Clearly this policy exactly satisfies the budget constraint (i.e., is a feasible solution to the original optimization problem), and by boundedness of outcomes
$V(\nu^{+}_{\theta})=V(\nu^{*}_{\theta})+\mathcal{O}\left(\left\lVert\theta\right\rVert_{2}/\varepsilon\right).$
Now, because $\nu^{+}_{\theta}$ is feasible we have
$\omega(0)\leq-V(\nu^{+}_{\theta})=-V(\nu^{*}_{\theta})+\mathcal{O}\left(\left\lVert\theta\right\rVert_{2}\right)=\omega(\theta)+\mathcal{O}\left(\left\lVert\theta\right\rVert_{2}\right),$
which implies continuity of $\omega$ at 0 because, by construction, $\omega$ is decreasing in $\theta$.
Appendix C Concentration Bounds
[…]
$\sup_{\left\lVert p-p^{*}\right\rVert\leq\delta_{n}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\left(F_{in}(p)-F_{in}(p^{*})-\left(f_{n}(p)-f_{n}(p^{*})\right)\right)\right\rvert=o_{p}\left(\frac{1}{\sqrt{n}}\right).$ (73)
The value in an asymptotic equicontinuity expansions is that we may often be able to assume that the expected function $f_{n}(p)$ is differentiable even when the $F_{in}(p)$ themselves are not (e.g., if the $F_{in}(p)$ capture choices regarding supply and demand that may vary discontinuously in prices). Asymptotic equicontinuity then allows us to approximate sample averages of $F_{in}(p)$ by Taylor expanding $f_{n}(p)$.
The goal of this section is to establish asymptotic equicontinuity for a number of function classes used throughout the proofs. Our argument relies on technical tools from empirical process theory that go beyond what’s used elsewhere in the paper, and here we will only give brief references to the results needed to establish our desired claims. Section 2 of van der Vaart & Wellner (1996) provides an excellent introduction and reference to the tools underlying the arguments made here.
Our approach to establishing asymptotic equicontinuity of relevant quantities in our marketplace model starts by bounding the bracketing number of approximately monotone functions. Let $F_{i}:\mathbb{R}^{J}\rightarrow\mathbb{R}$ be random functions belonging to some set $\mathcal{F}$, and let $\mathcal{S}\subset\mathbb{R}^{J}$ be the set of possible market equilibrium prices. For any pair $p_{-},\,p_{+}$ in $\mathbb{R}^{J}$, we define the $\mathcal{F}$-bracket over $\mathcal{S}$ as
$[p_{-},\,p_{+}]_{\mathcal{F}}=\left\{p\in\mathcal{S}:F(p_{-})\geq F(p)\geq F(p_{+})\text{ for all }F\in\mathcal{F}\right\}.$ (74)
Given a distribution $P$ over the random functions, i.e., $F_{i}\,{\buildrel\text{iid}\over{\sim}\,}P$, we define the $L_{2}(P)$-length of the bracket as $d^{2}_{P}(p_{-},\,d_{+})=\mathbb{E}_{P}\left[(F_{i}(p_{-})-F_{i}(p_{+}))^{2}\right]$, and the $\varepsilon$-bracketing number of $\mathcal{S}$ under $P$ as
$N_{[]}(\varepsilon,\,\mathcal{S},\,L_{2}(P))=\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}[p_{-}(k),\,p_{+}(k)]_{\mathcal{F}}\text{ with }d_{P}\left(p_{-}(k),\,p_{+}(k)\right)\leq\varepsilon\text{ for all }k\right\}.$ (75)
The following result bounds the bracketing number of functions that are approximately monotone and weakly continuous in the sense of Assumption 3.
Lemma 18.
Consider a class of random functions $F_{i}:\mathbb{R}^{J}\rightarrow\mathbb{R}$ with $F_{i}\in\mathcal{F}$ almost surely, and let $\mathcal{S}$ is a compact subset of $\mathbb{R}^{J}$. Suppose that, for all $p\in\mathbb{R}^{J}$, $\varepsilon\leq 1$ and $\left\lVert\delta\right\rVert_{2}\leq C\varepsilon$, we have
$F_{i}(p-\varepsilon e_{j})\geq F_{i}(p+\delta)\geq F_{i}(p+\varepsilon e_{j}).$ (76)
Suppose also that the $F_{i}$ have a distribution $P$ under which
$\mathbb{E}_{P}\left[\left(F_{i}(p)-F_{i}(p^{\prime})\right)^{2}\right]\leq L\left\lVert p-p^{\prime}\right\rVert_{2}\text{ for all }p,\,p^{\prime}\in\mathbb{R}^{J}.$ (77)
Then, there exists $\theta(\mathcal{S},\,C,\,L)$ depending only on the geometry of $\mathcal{S}$ and the constants given above such that, for all $\varepsilon\leq 1$,
$N_{[]}(\varepsilon,\,\mathcal{S},\,L_{2}(P))\leq\theta(\mathcal{S},\,C,\,L)\,\varepsilon^{-2J}.$ (78)
Proof.
Let $B_{\delta}(p)=\left\{p^{\prime}\in\mathcal{S}:\left\lVert p^{\prime}-p\right\rVert_{2}\leq\delta\right\}$. By (76), we have
$B_{C\zeta}(p)\subseteq[p-\zeta e_{j},\,p+\zeta e_{j}]_{\mathcal{F}}.$
Meanwhile, by (77) we have $d_{P}^{2}\left(p-\zeta e_{j},\,p+\zeta e_{j}\right)\leq 2L\zeta$, and so $d_{P}\left(p-\zeta e_{j},\,p+\zeta e_{j}\right)\leq\varepsilon$ if $\zeta=\varepsilon^{2}/(2L)$. These two facts together imply that
$N_{[]}(\varepsilon,\,\mathcal{S},\,L_{2}(P))\leq\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}B_{C(\varepsilon^{2}/(2L))}(p_{k})\text{ with }p_{k}\in\mathbb{R}^{J}\right\},$
where the right-hand side quantity is the $C(\varepsilon^{2}/(2L))$-covering number of $\mathcal{S}$ under the usual norm .$\left\lVert.\right\rVert_{2}$. For $\varepsilon\leq 1$, this quantity can be further be bounded as (Polyanskiy & Wu 2024, Theorem 27.3)
$\begin{split}&\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}B_{C(\varepsilon^{2}/(2L))}(p_{k})\text{ with }p_{k}\in\mathbb{R}^{J}\right\}\\ &\quad\quad\quad\quad\leq 3^{J}\left(C\left(\frac{\varepsilon^{2}}{2L}\right)\right)^{-J}\frac{\operatorname{vol}\left(\operatorname{conv}\left(\mathcal{S}*B_{C/(2L)}\right)\right)}{\operatorname{vol}\left(B_{1}\right)},\end{split}$ (79)
where $B_{r}$ denotes the radius-$r$ ball in $\mathbb{R}^{J}$ centered at 0, $\mathcal{A}*\mathcal{B}=\left\{a+b:a\in\mathcal{A},\,b\in\mathcal{B}\right\}$, $\operatorname{conv}(.)$ denotes the convex hull of a set and $\operatorname{vol}(.)$ denotes the volume of a set. The scaling in (79) establishes the desired claim. ∎
Lemma 19.
We say that a triangular array of random functions $F_{in}$ is $\zeta$-subgaussian over $\mathcal{S}$ if there exists a constant $C$ such that, for all $n$ and $t\geq 1$,
$\mathbb{P}\left[\sup_{p\in\mathcal{S}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}F_{in}(p)-f(p)\right\rvert>\frac{t}{\sqrt{n}}\right]\leq Ct^{\zeta}e^{-2t^{2}}.$ (80)
Under Assumptions 3, 4 and 5, let $U_{i}$ be random price perturbations with $\mathbb{E}\left[U_{i}\right]=0$ and $\left\lvert U_{i}\right\rvert\leq h_{n}$ almost surely for some sequence $h_{n}\rightarrow 0$. Suppose furthermore that $W_{i}$ is Bernoulli-randomized as in Design 1. Then, the following function classes are $\zeta$-subgaussian with $\zeta=2J$:
[…]
  • $Z_{i}(W_{i},\,p+U_{i})$ and $(W_{i}/\pi_{i}-(1-W_{i})/(1-\pi_{i}))Z_{i}(W_{i},\,p+U_{i})$.
[…]
  • $Y_{i}(W_{i},\,p+U_{i})$ and $(W_{i}/\pi_{i}-(1-W_{i})/(1-\pi_{i}))Y_{i}(W_{i},\,p+U_{i})$.
Proof.
The result for $Z_{i}(w,\,p+U_{i})$ and $Z_{i}(W_{i},\,p+U_{i})$ follows immediately from the bracketing bounds established in Lemma 18, boundedness as required in Assumption 3, and Theorem 2.14.9 of van der Vaart & Wellner (1996). The result for $W_{i}/\pi_{i}\,Z_{i}(W_{i},\,p+U_{i})$ and $(1-W_{i})/(1-\pi_{i})\,Z_{i}(W_{i},\,p+U_{i})$, and thus also that of their difference, follows by the same argument because $W_{i}/\pi_{i}$ and $(1-W_{i})/(1-\pi_{i})$ are positive and bounded by $\eta^{-1}$, and thus these functions satisfy the conditions of Lemma 18 with a constant $L/\eta$ in (77).
Next recall to functions constructed in therms of $Y_{i}(w,\,p)$, recall that by Assumption 3 there is a random function $H_{i}(w,\,p)$ that is Lipschitz continuous in $p$ such that
$Y_{i}(w,\,p+U_{i})$ $=H_{i}(w,\,p+U_{i})+\left(\alpha+\beta^{\top}(p+U_{i})\right)\cdot Z_{i}(w,\ p+U_{i})$ (81)
$=H_{i}(w,\,p+U_{i})+\alpha^{\top}Z_{i}(w,\ p+U_{i})+p^{\top}\beta\,Z_{i}(w,\ p+U_{i})+U_{i}^{\top}\beta Z_{i}(w,\ p+U_{i}).$
Our above concentration result immediately generalizes to the 2nd and 3rd terms above, because the multipliers are non-random and $p$ takes values within a compact set $\mathcal{S}$. To control the 4th term, we note that
$U_{i}^{\top}\beta Z_{i}(w,\ p+U_{i})=\sum_{j,\,k=1}^{J}\left(U_{ij}\beta_{jk}\right)_{+}Z_{ik}(w,\ p+U_{i})-\sum_{j,\,k=1}^{J}\left(-U_{ij}\beta_{jk}\right)_{+}Z_{ik}(w,\ p+U_{i}),$
i.e., the term can be written as $2J^{2}$ weighted averages of the net demand for a given good, with non-negative, bounded weights. We can thus obtain concentration for this term by applying our above argument a finite number of times.
Finally, to control the contribution of the $H_{i}(w,\,p)$ term, we note that bracketing bounds for Lipschitz function classes are standard; see, e.g., Section 2.7.4 of van der Vaart & Wellner (1996). Adapting these results to our setting, let
[…]
Clearly, for any $p\in\mathbb{R}^{J}$,
[…]
and $d_{P}((p,\,-M\varepsilon),\,(p,\,M\varepsilon))=2M\varepsilon$ for any distribution $P$ over the functions $A(\cdot)$. Thus, using notation from Lemma 18,
$N_{[]}(\varepsilon,\,\mathcal{S},\,L_{2}(P))\leq\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}B_{\varepsilon/(2M)}(p_{k})\text{ with }p_{k}\in\mathbb{R}^{J}\right\},$ (82)
which is in turn $O(\varepsilon^{-J})$ by the same argument as used in (79). This bracketing bound can be applied to $H_{i}(w,\,p+U_{i})$, $H_{i}(W_{i},\,p+U_{i})$, etc., thus enabling us to again apply Theorem 2.14.9 of van der Vaart & Wellner (1996). Other functions constructed using $Y_{i}(w,\,p)$ can be controlled analogously. ∎
Lemma 20.
Under the setting of Lemma 19, the following function classes are asymptotically equicontinuous, i.e., they satisfy (73) at all $p\in\mathcal{S}$:
[…]
  • $Z_{i}(W_{i},\,p+U_{i})$ and $(W_{i}/\pi_{i}-(1-W_{i})/(1-\pi_{i}))Z_{i}(W_{i},\,p+U_{i})$.
[…]
  • $Y_{i}(W_{i},\,p+U_{i})$ and $(W_{i}/\pi_{i}-(1-W_{i})/(1-\pi_{i}))Y_{i}(W_{i},\,p+U_{i})$.
Furthermore, the residuals in the asymptotic equicontinuity expansion,
$R_{n}=\sup\left\{\left\lvert\frac{1}{n}\sum_{i=1}^{n}\left(F_{in}(p)-F_{in}(p^{*})-\left(f_{n}(p)-f_{n}(p^{*})\right)\right)\right\rvert:\left\lVert p-p^{*}\right\rVert\leq\delta_{n},\ p\in\mathcal{S}\right\}$ (83)
satisfy $\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[R_{n}^{2}\right]=0$.
Proof.
We start with the functions constructed using $Z_{i}(w,\,p)$. Fix a good $j\in\left\{1,\,\ldots,\,J\right\}$. By Assumption 3, we see that $Z_{ij}(w,\,p+U_{i})$ and $Z_{ij}(W_{i},\,p+U_{i})$ satisfy the conditions of Lemma 18; and, by weak continuity of $Z_{ij}(w,\,p)$, we have
$\lim_{n\rightarrow\infty}\operatorname{Cov}\left[Z_{ij}(w,\,p+U_{i}),\ Z_{ij}(w,\,p^{\prime}+U_{i})\right]=\operatorname{Cov}\left[Z_{ij}(w,\,p),\ Z_{ij}(w,\,p^{\prime})\right].$ (84)
Thus, for both $F_{in}(p)=Z_{ij}(w,\,p+U_{i})$ and $F_{in}(p)=Z_{ij}(W_{i},\,p+U_{i})$, weak convergence of the empirical process
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(F_{in}(p)-f_{n}(p)\right)$
to a Gaussian process indexed by $p$ follows from Lemma 18 and Theorem 2.11.23 of van der Vaart & Wellner (1996). Furthermore, asymptotic equicontinuity follows by weak continuity as required in Assumption 5, as weak continuity implies that the limiting Gaussian process has continuous sample paths.
Asymptotic equicontinuity of $(W_{i}/\pi_{i}-(1-W_{i})/(1-\pi_{i}))Z_{ij}(W_{i},\,p+U_{i})$ follows analogously via the argument made in Lemma 19. We have thus shown asymptotic equicontinuity separately for the $j$-th entry of each function class constructed using $Z_{i}(w,\,p)$; and this implies asymptotic equicontinuity of the full (vector-valued) functions themselves also.
To verify the 2nd-moment bounds on the residuals in the expansions, we first note that
$R_{n}$ $=\sup_{\left\lVert p-p^{*}\right\rVert\leq\delta_{n}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\left(F_{in}(p)-F_{in}(p^{*})-\left(f_{n}(p)-f_{n}(p^{*})\right)\right)\right\rvert$
$\quad\quad\quad\quad\quad\quad\quad\quad\leq 2\sup_{p\in\mathcal{S}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\left(F_{in}(p)-f_{n}(p)\right)\right\rvert,$
which has a rapidly decaying tail by Lemma 19. Thus, in particular,
$\limsup_{n\rightarrow\infty}n^{2}\,\mathbb{E}\left[R_{n}^{4}\right]<\infty.$
Now, pick any $c>0$. By asymptotic equicontinuity, $\limsup_{n\rightarrow\infty}\mathbb{P}\left[\left\lvert R_{n}\right\rvert>c/\sqrt{n}\right]=0$. Furthermore
$\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[R_{n}^{2}\right]$ $=\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[1\left(\left\{\left\lvert R_{n}\right\rvert\leq c/\sqrt{n}\right\}\right)R_{n}^{2}\right]+n\,\mathbb{E}\left[1\left(\left\{\left\lvert R_{n}\right\rvert>c/\sqrt{n}\right\}\right)R_{n}^{2}\right]$
$\leq c^{2}+\limsup_{n\rightarrow\infty}\sqrt{\mathbb{P}\left[\left\lvert R_{n}\right\rvert>c/\sqrt{n}\right]}\sqrt{n^{2}\,\mathbb{E}\left[R_{n}^{4}\right]}$
$\leq c^{2}+\limsup_{n\rightarrow\infty}\sqrt{\mathbb{P}\left[\left\lvert R_{n}\right\rvert>c/\sqrt{n}\right]}\limsup_{n\rightarrow\infty}\sqrt{n^{2}\,\mathbb{E}\left[R_{n}^{4}\right]}$
$=c^{2},$
where for the second inequality we used Cauchy-Schwarz. Since this holds for any $c>0$, we see that in fact $\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[R_{n}^{2}\right]=0$.
To establish asymptotic equicontinuity of the functions constructed using $Y_{i}(w,\ p)$, we again leverage the decomposition (81), and verify that each term is asymptotically equicontinuous on its own. We again focus on deriving asymptotic equicontinuity for $F_{i}(p)=Y_{i}(w,\ p+U_{i})$; the extension to other functions is immediate. For the first term in (81), our Lipschitz assumption yields
$\lim_{n\rightarrow\infty}\operatorname{Cov}\left[H_{i}(w,\,p+U_{i}),\ H_{i}(w,\,p^{\prime}+U_{i})\right]=\operatorname{Cov}\left[H_{i}(w,\,p),\ H_{i}(w,\,p^{\prime})\right],$ (85)
and so we can use (82) and Theorem 2.11.23 of van der Vaart & Wellner (1996) to establish weak convergence to a Gaussian process; and continuity of the limiting process is immediate from Lipschitz continuity of $H_{i}(w,\,p)$. Meanwhile, asymptotic equicontinuity of the 2nd, 3rd and 4th terms follows by adapting arguments already used to study net demand, and discussed in the proof of Lemma 19. ∎
Appendix D Additional Results
This appendix is for results that are cited but not included in the main text.
D.1 Market-Clearing
Proposition 21.
Single-Good Market. In a market with $J=1$ goods, assume that outcomes and net demand functions are bounded, and for all $p\in\mathcal{S}$ and $w\in\{0,1\}$, $\mathbb{P}(Z_{i}(w,p)\text{ is continuous at }p)=1$ and $\mathbb{P}(Y_{i}(w,p)\text{ is continuous at }p)=1$. In addition, assume that
  • $\mathbb{E}[Z_{i}(w,0)]>0$ and there exists a finite $b>0$ such that $\mathbb{E}[Z_{i}(w,b)]<0$ for $w\in\{0,1\}$.
  • With probability 1, $Z_{i}(w,p)$ is monotonically non-increasing in $p$ for $w\in\{0,1\}$.
Let $Z_{n}(\bm{w},p)=\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(w_{i},\bm{p})$ and $P_{n}(\bm{w})=\arg\min\limits_{p}|Z_{n}(\bm{w},p)|$ for $\bm{w}\in\{0,1\}^{n}$. In cases where there are multiple minimizers, choose one by some deterministic rule.
[…]
$\mathcal{S}_{\mathbf{w}}=\left\{p\in\mathbb{R}^{J}:\left\lVert\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(w_{i},p)\right\rVert_{2}\leq a_{n}\right\}.$ (86)
is non-empty with probability $1-e^{-cn}$. And $P_{n}(\bm{w})\in\mathcal{S}_{\mathbf{w}}$ when it is non-empty.
Proof.
Denote by $\mathcal{E}_{n}$ the event that the realized $Z_{n}(\bm{w},p)$ crosses zero during the interval $(0-s,b+s)$:
$\mathcal{E}_{n}=\{Z_{n}(0)>0,\mbox{ and }Z_{n}(b)<0\},$ (87)
and denote by $E_{n}$ the indicator variable for $\mathcal{E}_{n}$: $E_{n}=\mathbb{I}(\mathcal{E}_{n})$.
Suppose that $E_{n}=1$. Here, either $Z_{n}(P_{n})=0$, or $|Z_{n}(P_{n})|\neq 0$, and we will focus on the latter case. Suppose $|Z_{n}(P_{n})|\neq 0$. Because we condition on $E_{n}=1$, we have that $Z_{n}(0)>0$ and $Z_{n}(b)<0$. Since $P_{n}$ is chosen to be a minimizer of $|Z_{n}(p)|$, the only possibility is that the function $Z_{n}(\cdot)$ crosses zero at a point of discontinuity, $P^{\prime}$, that is, $Z_{n}(\tilde{p})<0$ for all $\tilde{p}>P^{\prime}$, and $Z_{n}(\bar{p})>0$ for all $\bar{p}<P^{\prime}$.
We have assumed that $\mathbb{P}(Z_{i}(W_{i},\cdot)\mbox{ is discontinuous at $p$})=0$ for any $p>0$, the $Z_{i}$’s are drawn independently, and by the boundedness assumption they are bounded in magnitude by $M$. These facts further imply that with probability one, any jump in $Z_{n}(\cdot)$ cannot exceed magnitude $M/n$, for otherwise it would have required at least two separate $Z_{i}$ to have a point of discontinuity at the same exact location. Formally, conditional on $E_{n}=1$, we have that with probability one:
$|Z_{n}(P_{n})|\leq|\lim_{p\uparrow P^{\prime}}Z_{n}(p)-\lim_{p\downarrow P^{\prime}}Z_{n}(p)|\leq M/n,$ (88)
Choosing $a_{n}=M/n$, we note that $\lim\limits_{n\to\infty}a_{n}\sqrt{n}=0$, and that we have that the specified $P_{n}(\bm{w})\in\mathcal{S}_{\mathbf{w}}$ whenever $\mathcal{S}_{\mathbf{w}}$ is non-empty. It is non-empty whenever $E_{n}=1$, so to finish the proof we now check the probability that $E_{n}=1$.
If $E_{n}=0$, there are four possible cases. The first two cases are $Z_{n}(0)=0$ or $Z_{n}(b)=0$, in which case $Z_{n}(P_{n})=0$. The last two are $Z_{n}(p)>0$ for all $p\in\mathcal{S}$, or $Z_{n}(p)<0$ for all $p\in\mathcal{S}$, in which case $Z_{n}(P_{n})\leq c_{1}$ for some constant $c_{1}>0$, since each net demand function is bounded. Note that $c_{1}$ is not less than the specified $a_{n}$, so to verify the proposition we need that $\mathbb{P}(E_{n}=0)$ is exponentially small. To bound $\mathbb{P}(E_{n}=0)$,
$\mathbb{P}(E_{n}=0)$ $\leq\mathbb{P}(Z_{n}(0)\leq 0)+\mathbb{P}(Z_{n}(b)\geq 0)$
$=\mathbb{P}(Z_{n}(0)-\mathbb{E}[Z_{n}(0)]\leq-\mathbb{E}[Z_{n}(0)])+\mathbb{P}(Z_{n}(b)-\mathbb{E}[Z_{n}(b)]\geq-\mathbb{E}[Z_{n}(b)])$
$\leq\mathbb{P}(|Z_{n}(0)-\mathbb{E}[Z_{n}(0)]|\geq\mathbb{E}[Z_{n}(0)])+\mathbb{P}(|Z_{n}(b)-\mathbb{E}[Z_{n}(b)]|\geq-\mathbb{E}[Z_{n}(b)])$
$\leq\mathbb{P}(|Z_{n}(0)-\mathbb{E}[Z_{n}(0)]|\geq\varepsilon)+\mathbb{P}(|Z_{n}(b)-\mathbb{E}[Z_{n}(b)]|\geq\varepsilon),$
for $\varepsilon=\min\{\mathbb{E}[Z_{n}(0),-\mathbb{E}[Z_{n}(c)]\}>0$. We can now use Hoeffding’s inequality, and obtain that for any $n$ and $\varepsilon>0$,
$\mathbb{P}(E_{n}=0)$ $\leq 2\cdot\exp\left(\frac{-2n\varepsilon^{2}}{4c_{1}^{2}}\right),$
$=2\cdot\exp\left(\frac{-n\varepsilon^{2}}{2c_{1}^{2}}\right).$ (89)
We have now verified that $\mathcal{S}_{\mathbf{w}}$ is non-empty with probability $1-e^{-cn}$, for some $c>0$, which completes the proof.
D.2 Convergence Rate of $\tau_{\text{AIE}}$ to $\tau^{*}_{\text{AIE}}$
In this section, we provide a simple example where $Z_{i}(w,p)$ and $Y_{i}(w,p)$ are differentiable in $p$. In this example, the asymptotic representation of $\tau_{\text{AIE}}-\tau^{*}_{\text{AIE}}$ depends on $\frac{1}{n}\sum\limits_{i=1}^{n}\nabla_{p}Y_{i}(w,p)-y(w,p)$, and it is straightforward to extend this example so that it also depends on the concentration of $\frac{1}{n}\sum\limits_{i=1}^{n}\nabla_{p}Z_{i}(w,p)$ around its expected derivative. This suggests that it is not possible to get $\sqrt{n}$ converge of $\tau_{\text{AIE}}$ to $\tau^{*}_{\text{AIE}}$ under Assumption 5, which allows for discontinuous individual-level demand and outcome functions.
Define bounded random variables $(Y_{i}(1),Y_{i}(0),Z_{i}(1),\beta_{i})$ that are sampled IID from some distribution. $Z_{i}(0)=0$ always. This distribution is such that $\mathbb{E}[Z_{i}(1)]=z(1)$, $\mathbb{E}[Y_{i}(1)-Y_{i}(0)]=y(1)-y(0)]$ and $\mathbb{E}[\beta_{i}]=\beta$. $W_{i}$ is drawn IID from $\mbox{Bernoulli}(\pi)$. The rest of the data-generating process is:
$\begin{split}Z_{i}(w,p)=\theta_{0}-p+Z_{i}(W_{i}),\qquad Y_{i}(w,p)=-\beta_{i}W_{i}p+Y_{i}(W_{i}).\end{split}$
In this model, $p^{*}_{\pi}=\theta_{0}+\pi\cdot z(1)$. The population AIE is
$\tau^{*}_{\text{AIE}}=-\pi\cdot\beta\cdot z(1).$
This model implies that $P_{n}(W_{i}=1;\bm{W}_{-i})-P_{n}(W_{i}=0;\bm{W}_{-i})=\frac{1}{n}Z_{i}(1)$ and $Y_{j}(W_{j},P_{n}(W_{j}=1;\bm{W}_{-i})-Y_{j}(W_{j},P_{n}(W_{j}=0;\bm{W}_{-i})=-\beta_{j}W_{j}\cdot\frac{1}{n}Z_{i}(1)$.
This implies that $\tau_{\text{AIE}}$ in this model is:
$\begin{split}\tau_{\text{AIE}}&=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{j\neq i}\pi\cdot\beta_{j}\frac{1}{n}(Z_{i}(1)-Z_{i}(0))\\ &=\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(1)\frac{1}{n}\sum\limits_{j=1}^{n}\pi\cdot\beta_{j}\\ &=\tau^{*}_{\text{AIE}}+\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(1)\frac{1}{n}\sum\limits_{j=1}^{n}\pi\cdot(\beta_{j}-\beta)+\frac{1}{n}\sum\limits_{i=1}^{n}(Z_{i}(1)-Z_{i}(0))-z(1)-z(0))\pi\cdot\beta\\ &\stackrel{{\scriptstyle(1)}}{{=}}\tau^{*}_{\text{AIE}}+z(1)\frac{1}{n}\sum\limits_{j=1}^{n}\pi\cdot(\beta_{j}-\beta)+\frac{1}{n}\sum\limits_{i=1}^{n}(Z_{i}(1)-z(1))\pi\cdot\beta+o_{p}(n^{-1/2}).\end{split}$
The last line is by the application of the LLN and recognizing that the CLT applies to the term $\frac{1}{n}\sum\limits_{j=1}^{n}\pi\cdot(\beta_{j}-\beta)$. In this case there is $\sqrt{n}$ convergence of $\tau_{\text{AIE}}$ to $\tau^{*}_{\text{AIE}}$, but the distribution depends on the variance of $\nabla_{p}Y_{j}(1,p^{*}_{\pi})$. In a more general derivation with non-differentiable $Y_{i}$, the analogous term would not converge at a $\sqrt{n}$ rate, so the scaling of $\tau_{\text{AIE}}$ to $\tau^{*}_{\text{AIE}}$ will be slower than $\sqrt{n}$ under the assumptions in the paper (just as the convergence of $\hat{\tau}_{\text{AIE}}$ to $\tau^{*}_{\text{AIE}}$ is slower than $\sqrt{n}$).
D.3 Observing a Sub-Sample of the Market
The estimators for $\tau^{*}_{\text{ADE}}$ and $\tau^{*}_{\text{AIE}}$ are unchanged when a sub-sample of the $n$ agents in the finite-sample market are observed rather than all of them. However, the expansion for $\hat{\tau}_{\text{ADE}}$ described in Theorem 5 is affected slightly, which changes the asymptotic variance. This appendix provides some detail on these claims which were made in the main text, under the assumptions of Theorem 5.
Let $n$ be the size of the finite-sample market. Let $\mathcal{M}\subset\{1,\ldots,n\}$ index the $m$ total number of observations sampled without replacement from the finite-sample market. We assume that the market price clears the finite-sample market, but we only observe $Y_{j}$ and $Z_{j}$ for $j\in\mathcal{M}$. Let $\hat{s}=\frac{1}{m}\sum\limits_{j\in M}W_{j}$ be the fraction of individuals treated in the sub-sample, $\hat{t}=\frac{m}{n}$ be the fraction of individuals observed, and $\hat{\pi}=\frac{1}{n}\sum\limits_{i=1}^{n}W_{i}$ be the fraction of individuals treated in the entire finite market. We assume that for all agents that are not in $\mathcal{M}$, then $W_{i}=0$. The sub-sample is a fixed fraction of the finite market, so that as $n\to\infty$, $\hat{s}\to s>0$, $\hat{\pi}\to\pi>0$ and $\hat{t}\to t>0$.
The estimators $\hat{\tau}_{\text{ADE}}$ and $\hat{\tau}_{\text{AIE}}$ are exactly the same as in the main text, except the observations used are in $\mathcal{M}$ rather than in $\{1,\ldots n\}$. Let $h_{n}$ remain the same as in the main text, where the shrinkage rate of the price perturbations depend on $n$.
$\begin{split}&\hat{\tau}_{\text{m, ADE}}=\frac{1}{m}\sum\limits_{j\in\mathcal{M}}\left[\frac{W_{j}Y_{j}}{\hat{s}}-\frac{(1-W_{j})Y_{j}}{1-\hat{s}}\right]\\ &\hat{\tau}_{\text{m, AIE}}=-\hat{\gamma}^{\top}\frac{1}{m}\sum\limits_{j\in\mathcal{M}}\left[\frac{W_{j}{Z}_{j}}{\hat{s}}-\frac{(1-W_{j}){Z}_{j}}{1-\hat{s}}\right].\end{split}$
Let $\bm{Y}$ be the $m$-length vector of observed outcomes, $\bm{U}$ is the $m\times J$ matrix of observed price perturbations, and $\bm{Z}$ is the $m\times J$ matrix of observed net demand. Then, $\hat{\gamma}=(\bm{U}^{\top}\bm{Z})(\bm{U}^{\top}\bm{Y})$.
The asymptotic expansion for $\hat{\tau}_{m,\text{AIE}}$ is unchanged, apart from depending on $m$ rather than $n$, since the variance of $P_{n}-p^{*}_{\pi}$ does not impact the limiting distribution.
$\hat{\tau}_{m,\text{AIE}}=\tau^{*}_{\text{AIE}}-Q_{z}^{\top}\frac{1}{\sqrt{m}h^{2}_{n}}\sum\limits_{j\in\mathcal{M}}U_{j}\nu_{j}(W_{j})+o_{p}(1),$
where $\nu_{j}(W_{j})=Y_{j}(W_{j},p^{*}_{\pi})-Z_{j}(W_{j},p^{*}_{\pi})^{\top}[\xi_{z}^{-1}]^{\top}\xi_{y}$ and $Q_{z}=\xi_{z}^{-1}\tau^{*,z}_{\text{ADE}}$. The asymptotic variance can still be estimated by $\hat{\sigma}_{I}$, where each component is estimated using the observed data only.
The asymptotic expansion for $\hat{\tau}_{\text{ADE}}$ is affected, since $P_{n}-p^{*}_{\pi}$ does impact the asymptotic expansion, and the finite sample market price depends on $n$ rather than $m$.
$\begin{split}\hat{\tau}_{m,\text{ADE}}&\stackrel{{\scriptstyle(1)}}{{=}}\tau^{*}_{\text{ADE}}+\frac{1}{n}\sum\limits_{i=1}^{n}\left[\frac{\mathbbm{1}(i\in\mathcal{M})}{\hat{t}}\left(\frac{W_{i}\varepsilon_{i}(1)}{\hat{s}}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\hat{s}}\right)-\Delta_{i}(W_{i},p^{*}_{\pi})\right]+o_{p}(n^{-0.5})\\ &=\tau^{*}_{\text{ADE}}+\frac{1}{n}\sum\limits_{i=1}^{n}\left[\frac{\mathbbm{1}(i\in\mathcal{M})}{t}\left(\frac{W_{i}\varepsilon_{i}(1)}{s}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-s}\right)-\Delta_{i}(W_{i},p^{*}_{\pi})\right]+o_{p}(n^{-0.5})\\ &\stackrel{{\scriptstyle(2)}}{{=}}\tau^{*}_{\text{ADE}}+\frac{1}{n}\sum\limits_{i=1}^{n}\frac{\mathbbm{1}(i\in\mathcal{M})}{t}\left(\frac{W_{i}\varepsilon_{i}(1)}{s}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-s}-t\Delta_{i}(W_{i},p^{*}_{\pi})\right)\\ &\qquad-\frac{1}{n}\sum\limits_{i=1}^{n}\mathbbm{1}{(i\notin\mathcal{M})}\Delta_{i}(W_{i},p^{*}_{\pi})+o_{p}(n^{-0.5})\end{split}$
(1) comes from following the same steps as in the proof of Theorem 5. (2) comes from splitting up the sum for observations in $\mathcal{M}$ and those not in $\mathcal{M}$. As before $\varepsilon_{i}(w)=Y_{i}(w,p^{*}_{\pi})-y(w,p^{*}_{\pi})$.
The CLT now applies to this expansion:
$\begin{split}&\sqrt{n}(\hat{\tau}_{m,\text{ADE}}-\tau^{*}_{\text{ADE}})\Rightarrow\mathcal{N}(0,\sigma^{2}_{t,D})\\ &\sigma^{2}_{t,D}=\frac{1}{t}\mathbb{E}\left[\left(\frac{W_{i}\varepsilon_{i}(1)}{s}-\frac{(1-W_{i})\varepsilon_{i}(0)}{(1-s)}-t\Delta_{i}(W_{i},p^{*}_{\pi})\right)^{2}\right]+(1-t)\mathbb{E}\left[\Delta_{i}(W_{i},p^{*}_{\pi})^{2}\right]\end{split}$
For inference on $\tau^{*}_{\text{ADE}}$ using the sub-sample only, we can estimate each component of $\sigma^{2}_{t,D}$ using only observations in $\mathcal{M}$.
D.4 Alternative Estimator of AIE Variance
In analyzing the asymptotic properties of $\sigma^{2}_{I}$ in the proof of Theorem 7 in Appendix B.5, we dropped a term that was asymptotically negligible, which was $\sqrt{n}h_{n}({\gamma}^{\top}(\hat{\tau}^{z}_{\text{ADE}}-\tau^{*,z}_{\text{ADE}})).$
$\tilde{\sigma}^{2}_{I}$ adds a plug-in estimator for the second term, including an estimator for the variance of $\hat{\tau}^{z}_{\text{ADE}}$. $\tilde{\sigma}^{2}_{I}=\hat{\sigma}^{2}_{I}+h^{2}_{n}\hat{\gamma}^{\top}\hat{\sigma}^{2}_{z,D}\hat{\bm{\gamma}},$ where $\hat{\sigma}^{2}_{z,D}=\frac{1}{n}\sum\limits_{i=1}^{n}B_{i}B_{i}^{\prime}$ and the $J\times 1$ vector $\hat{B}_{i}=\frac{W_{i}\hat{\varepsilon}^{z}_{i}(1)}{\hat{\pi}}-\frac{(1-W_{i})\hat{\varepsilon}^{z}_{i}(0)}{1-\hat{\pi}}-[\hat{\xi}_{z1}-\hat{\xi}_{z0}]^{\top}\hat{\xi}_{z}^{-1}{Z}_{i}$. ${\varepsilon}^{z}_{i}(1)=Z_{i}-\frac{1}{n_{w}}\sum\limits_{i:W_{i}=w}Z_{i}$. Last, the $J\times J$ matrix $\hat{\xi}_{zw}$ for $w\in\{0,1\}$ are estimated from regressions of $Z_{i}$ on $U_{i}$ using only observations such that $W_{i}=w$. We found in simulations that this second-order correction led to better coverage properties at smaller sample sizes.